### v1

# Synthesis and stability of biomolecules under Earth’s upper mantle conditions

Abstract
How life started on Earth is a long-time unsolved mystery. There are various hypotheses ranging from outer space to seabed. Here, we applied extensive ab initio molecular dynamics (AIMD) simulations to study chemical reactions of NH3, H2O, H2, and CO at pressures (P) and temperatures (T) approximating the conditions of Earth’s upper mantle (i.e. 10 - 13 GPa, 1000 - 1400 K). Contrary to the previous assumption that larger organic molecules might readily dissociate in aqueous solutions at extreme P-T conditions, we found that many organic compounds formed and persisted in C-H-O-N fluids under these extreme conditions, including glycine, ribose, and uracil-like molecules. Particularly, our free energy calculations showed that the C-N bond is thermodynamically stable at 10 GPa and 1400 K. Moreover, our findings support the “RNA world” hypothesis, as we observed the exclusive formation of the 5-membered-ring form of ribose. By exploring the depths of Earth’s interior, we have uncovered a previously unexplored pathway through which life may have originated. These findings have contributed to our evolving understanding of the fundamental conditions necessary for life to arise on our planet.
I Introduction
The origin of life on Earth remains a profound scientific mystery. Numerous hypotheses have been put forward, but none have been conclusively verified perez2020prebiotic . As early as $\sim$150 years ago, Darwin proposed a “warm little pond” idea darwin1871jd , which was later developed by Oparin and Haldane into the influential ”primordial soup” theory in the 1920s oparin1957origin , that is, the reactions of small molecules, such as CH4, NH3, H2O, and CO2, leads to the abiotic origin of life on early Earth. In 1953, the famous Miller-Urey experiment was conducted to test this hypothesis at the presence of electric discharge miller1953production . Inspired by the ”primordial soup” theory, various prebiotic conditions have been proposed for the origin of life, ranging from the interstellar medium (ISM) herbst1995chemistry or other planets burton2012understanding ; sleep1998refugia to deep-sea hydrothermal vents takai2006ultramafics .
Hydrothermal vents are fissures in Earth’s oceanic crust where hot water rich in gases and minerals is released. Those deep sea vents are considered to have suitable conditions for origin of life. The pressure (P) of hydrothermal vents is tens of MPa depending on the depth of water, and the temperature (T) can reach 700 K. It is widely believed that the origin of life is unlikely to occur in the deeper regions of the Earth, such as the upper mantle, where the pressure can reach up to $\sim$13 GPa, accompanied by temperatures of $\sim$1700 K; the extremely high P-T conditions could readily degrade crucial biomolecules in aqueous solutions miller1988submarine . Otake et al. studied the stability of amino acids and their oligomerization at 10–5.0 GPa and 453–673 K, and found that high pressure inhibited the decomposition of amino acids ; however, their experiments were conducted at the water-poor condition otake2011stability . Previously, the fluids in Earth’s mantle were commonly modelled as simple mixtures of small volatile molecules, including H2O, CO2, CO, CH4, and H2 zhang2009model . However, recent experimental and theoretical studies have revealed the significant roles played by chemical speciation, aqueous ions, and complexes in the supercritical geofluids within Earth’s lithosphere manning2013chemistry ; sverjensky2020changing ; pan2016fate ; stolte2019large ; stolte2021water . These findings have challenged the previous assumptions and shed light on the potential existence of larger organic molecules.
It is very challenging to study the origin of life in laboratory settings due to the extreme environments involved. Atomistic simulations based on first principles have been applied to study prebiotic chemistry, offering significant molecular insights perez2020prebiotic . Saitta et al. applied ab initio molecule dynamics (AIMD) and metadynamics simulations to study the Miller-Urey experiment, and found unexpected reaction intermediates including formic acid and formamide saitta2014miller . Goldman et al. applied AIMD simulations to investigate the formation of prebiotic compounds from impact-induced shock compression of cometary ices followed by expansion to ambient conditions, and found that the impact of cometary ices could produce amino acids independently of atmospheric conditions and Earth minerals goldman2010synthesis ; goldman2013prebiotic ; koziol2015prebiotic . AIMD simulations coupled with enhanced sampling techniques have also been applied to study the role of mineral surfaces in prebiotic chemistry nair2006glycine ; pollet2006ab ; schreiner2011peptide . Despite the numerous previous studies, the exploration of prebiotic chemistry within Earth’s upper mantle remains an uncharted territory.
Here, we applied extensive AIMD simulations to study the C-H-O-N fluids at 10–13 GPa and 1000–1400 K, the P-T conditions as found at the bottom of Earth’s upper mantle. We discovered a large range of organic molecules directly formed without any catalysts at extreme P-T conditions, including glycine, ribose, and uracil-like molecules. Our free energy calculation suggested that the C-N bond is thermodynamically stable at 10 GPa and 1400 K. Our study has provided valuable insights into a new potential avenue for the origin of life within Earth’s interior, expanding our understanding of the conditions necessary for life to emerge.
II Results and Discussion
We initially mixed the H2O, H2, CO, and NH3 molecules in the simulation box. The molecular number of each molecule is 15. We conducted AIMD simulations at three P-T conditions: 10 GPa and 1000 K, 10 GPa and 1400 K, and 13 GPa and 1400 K to see how chemical products vary with P and T. Those P-T conditions are typically found at the bottom of Earth’s upper mantle. After 400 ps simulations at each P-T condition, we found about 100 different organic species. Here we mainly focus on the formation of bio-relevant molecules, so we classify the organic molecules into six categories: “CN”, “C2N”, “CN2”, “C2N2”, “C3Nx”, and “C4Nx”, based on the numbers of carbon and nitrogen atoms, as shown in Table 1. Fig. 1 shows the chemical structures of the representative species. Species with more than five carbon atoms are excluded due to their high instability and extremely short lifetime, typically lasting for only a few femtoseconds. The number and percentage of these CN-containing species as functions of time under three different P-T conditions are shown in Figures S5 and S6, respectively, which suggest that the concentrations of these species gradually stabilize over time.
At 10 GPa and 1000 K, most of the organic products belong to three main categories, as shown in Fig. 2(a). The “C2N2” species account for more than half of the CN-containing species, followed by the “CN” species, and finally “C2N”. Formamide is the primary product in the “CN” species saladino2012formamide . It contains the peptide bond, which is a significant precursor for the production of more complex biomolecules, such as amino acids pietrucci2015formamide , nucleic acids, proteins, and even sugars saladino2012genetics ; halfen2011formation . The “C2N” species has the same CN-backbone (“-N-C-C-”) as that of glycine, suggesting that the “C2N” species may be a potential reactant for the formation of amino acids. At this P-T condition, some of the “C2N” species are not stable and can be transformed into the “C2N2” species by reacting with ammonia or ammonium ions, so the CN-backbone of “C2N2” is “-N-C-C-N-”, and one interesting molecule was identified as alpha-hydroxy-glycineamide ($\alpha$-HGA).
With increasing the temperature to 1400 K along the isobar at 10 GPa, all the six categories of organic products at this P-T condition are observed. The “CN2” and “C4Nx” fractions are very small, while the other four species have similar fractions. Compared with 10 GPa and 1000 K, the portion of “C2N2” becomes smaller, while there are more “C2N” species. The “C2N” species formed here also share the same CN backbone (“-N-C-C-”) with glycine. Notably, in the “C2N” species, it is possible to form $\alpha$-hydroxyglycine, which can subsequently evolve into glycine through simple chemical reactions saitta2014miller . The “C2N2” species have the CN-backbone of“-N-C-C-N-”, including $\alpha$-HGA and oxamide. They are important organic molecules in the formation of the building blocks of peptides ligterink2019formation . Among the “CN” species, formamide remains the most abundant one. Additionally, the “CN2” species with the “-N-C-N-” backbone was observed, which did not typically form at 10 GPa and 1000 K. One representative product of this species is methanediamine (CH6N2) carballeira2001role , which was reported as a vital intermediate or precursor in the abiotic formation of nucleobases marks2022preparation .
With increasing the pressure to 13 GPa at 1400 K, the “CN” species becomes dominating, and the “C2N2” species comes next. For the “CN” species, in addition to formamide, some other crucial precursors for producing more complex biomolecules are identified, such as carbamic acid tossell2010happens , isocyanic acid belson1982preparation , and formimidic acid pranata1995computational . The “C2N2” species exhibit a distinct CN-backbone compared to the two systems at 10 GPa, forming as “-C-N-C-N-”, which is present in the organic matter of allophanic acid. This kind of acid plays a vital role in the synthesis of biomolecules (e.g. uracil), and would be used in subsequent simulations. Surprisingly, the fraction of “C2N” species sharply decreases to a small value, adopting a new CN-backbone structure of “-C-N-C-”, as observed in N-formylformamide sanz2009effects . This structural change makes it challenging for “C2N” species formed here to evolve into glycine-like molecules or larger amino acids. In the “CN2” species, we have discovered the presence of urea molecules. Urea is closely related with biological processes, so it would also serve as a valuable reactant for investigating the reactions to larger biomolecules.
For all the three P-T conditions, a common feature is that the “CN” and “C2N2” species are always the major compositions among the six CN-containing species. Fig. 2 (b) shows the lifetime of these CN-containing species. The lifetime distribution in Fig. 2 (b) shows that the “CN” and “C2N2” species could exist for a longer time at all the three P-T conditions, suggesting that they are relatively more stable than other species.
Among these generated chemical products, formamide is a crucial precursor molecule that plays a vital role in the synthesis of biomolecules in prebiotic chemistry saladino2012formamide ; pietrucci2015formamide ; enchev2021chemical ; saladino2015meteorite ; rios2015impact ; ferus2015high . Thus, we first investigated the formation of formamide. Fig. 3 (a) provides an illustration of the distinct formation pathways of formamide at three P-T conditions. The complete formation processes are shown in Figure S7. At 10 GPa and 1000 K, the formation process began with the dissociation of water molecules, generating OH- and H+ ions in the aqueous solution. When the H+ ion encountered a CO molecule, a reaction occurred, resulting in the formation of a C-H bond and an aldehyde group (-CHO). Subsequently, the aldehyde group reacted with the NH3 molecule, leading to the formation of a C-N bond, and ultimately, one N-H bond broke to produce formamide. At 10 GPa and 1400 K, the formation process differs from that at 10 GPa and 1000 K. One N-H bond in NH3 first broke, leading to the formation of a dissociative amino group (-NH2). The amino group then reacted directly with a CO molecule, forming the C-N bond. Later, the water molecule dissociated again, providing a hydrogen atom that ultimately contributed to the formation of the formamide molecule. At 13 GPa and 1400 K, the C-N bond was initially formed through the reaction of NH3 and CO molecules. After a very short time, the dissociated hydrogen atom from the water molecule also bonded to the carbon atoms, forming a C-H bond. Finally, by releasing a hydrogen atom from the NH3 group, the formamide molecule was generated. Although the formation pathways differ at three P-T conditions, we observed two common features: first, the reactants always include CO, H2O, and NH3, and second, the reaction pathways must involve the formation of C-N and C-H bonds and the breakage of N-H bonds.
For the generation of biomolecules in living matter, the formation of C-N bonds is the key step. We further applied the advanced sampling method (see details in method section) to calculate the free energy profile of the C-N bond formation at three P-T conditions, as shown in Fig. 3(b). The collective variable is the distance between the carbon and nitrogen atoms. At 10 GPa, 1000 K and 13 GPa, 1400 K, the C-N bonds, whose lengths are 1.33 Å and 1.31 Å  respectively, are at the metastable states. The free energy decreases upon dissociation of the C-N bond, indicating that breaking C-N bonds at these two P-T conditions are more favorable thermodynamically. The free energy profiles suggest that the energy barriers of the C-N bond dissociation are 17.57 kcal/mol and 16.40 kcal/mol, at 10 GPa, 1000 K and 13 GPa, 1400 K, respectively. As a comparison, at 10 GPa and 1400 K, the free energy increases upon dissociation of the C-N bond (1.31 Å), and the energy barrier is 59.04 kcal/mol, indicating that the C-N bond is thermodynamically and kinetically stable at this P-T condition. The free energy result agrees with the data in Fig 2, where there are more CN-containing species at 10 GPa and 1400 K than at the other two P-T conditions. Additionally, compared with the free energy profile at 10 GPa and 1000 K, there are two more energy minima at the C-N distance of 2.37 Å  and 2.45 Å at 10 GPa, 1400 K and 13 GPa, 1400 K, respectively. The free energy increases upon the dissociation of the C-N pair, indicating that these two minima are thermodynamically stable. The C-N distances are larger than that of a typical C-N covalent bond, which strongly implies the presence of an additional atom connecting the two. This is consistent with a wider variety of organic species at these two P-T conditions. Our free energy calculations suggest that the C-N bond formation is favorable at 10 GPa and 1400 K. With increasing the pressure to 13 GPa and keeping the temperature at 1400 K, the carbon and nitrogen atoms tend to be bonded together by a third atom.
After studying the basic chemical products at the three P-T conditions, we now turned to investigate the formation of large biomolecules as potential building blocks of life on Earth. Previous simulations have extensively studied glycine, the smallest amino acid, under various conditions saitta2014miller ; goldman2010synthesis ; wang2014discovering . In our simulations at 10 GPa and 1400 K, we discovered one molecule very similar to glycine, called $\alpha$-hydroxyglycine. Fig. 4 (a) illustrates the process of forming $\alpha$-hydroxyglycine from formamide. The chemical evolution suggests that, a hydrogen atom attacked the formamide molecule, leading to the creation of a protonated formamide molecule, also known as an aminohydroxymethyl radical. Subsequently, a CO2 molecule reacted with the protonated formamide, resulting in the production of a molecule with the “-N-C-C-” CN-backbone through forming a C-C bond. Finally, another hydrogen atom was captured by a “-COO” group, ultimately yielding the $\alpha$-hydroxyglycine molecule. In comparison to glycine, $\alpha$-hydroxyglycine differs by having one H atom and an “-OH” group instead of two H atoms connecting to the $\alpha$-C. Glycine can be obtained if the “-OH” group is replaced by a single H atom. During this process, the formation of the C-H bond is a crucial step. To identify appropriate P-T conditions for further advancing the chemical evolution of glycine, we calculated the free energy as a function of the C-H distance in Fig. 4 (b). Considering that the CN-backbone of the “C2N” species in the 13 GPa - 1400 K is “-C-N-C-”, very different from the CN-backbone in the glycine molecule, we did not consider this system. From the energy curves, we can see that the energy minima occurs at 1.07 Å for both 10 GPa - 1000 K and 10 GPa - 1400 K, and the energy barriers for C-H breaking are 21.91 kcal/mol and 22.04 kcal/mol, respectively, which are quite similar. However, the change of free energy from C-H bond formation to C-H bond dissociation is more positive at 10 GPa - 1400 K (6.76 kcal/mol) compared with that of 10 GPa - 1000 K (2.63 kcal/mol), which illustrates that 10 GPa - 1400 K condition is still the favorable one for forming C-H bond.
To test if we can also obtain glycine at extreme P-T conditions, we set up a new simulation box containing 5 $\alpha$-hydroxyglycine, 10 H2, and 10 H2O molecules and perform AIMD simulations at 10 GPa and 1400 K, as shown in Fig. 4(c). Initially, two $\alpha$-hydroxyglycine molecules would connect to a larger molecule, but later transformed into two new molecules after about 28 ps, dehydroglycine and glyoxylic acid. We mainly focus on dehydroglycine. As the simulation went on, this molecule possessed the amino group when one free proton was bonded to its N atom. After a shorter time, the $\alpha$-C also captured one proton, resulting in the final molecular structure of glycine at about 38 ps. Additionally, our observations indicate that once glycine is formed at 10 GPa - 1400 K, it remains at a stable state.
Both proteins and DNA are widely recognized as the important biomolecules for building blocks of life. However, the question of which one came first in the emergence of life remains a topic of debate. This is due to the fact that the replication of DNA relies on the catalytic activity of enzymes, which are a type of protein. Conversely, the production of proteins also needs the guidance of DNA, resulting in a chicken-and-egg dilemma. To solve this problem, a famous hypothesis known as “The RNA world” was proposed, which believed that RNA may occur earlier than DNA and protein gilbert1986origin ; nisbet1986origin . During this stage, the RNA not only have the ability to reproduce themselves and carry genetic information, but also could catalyze certain biological reactions in life activities like enzymemcclain1987model ; forster1990external . So the synthesis of RNA plays an important role in the early stage of the origins of life. Generally, RNA is composed of ribose sugar, four kinds of nucleobases, and a phosphate group. Considering that our simulation systems only contain C, H, O, and N elements, we then also explore whether the ribose and nucleobases could be synthesized under extreme conditions.
We first consider Ribose (C5H10O5), composed solely of C, H, O elements. The formose reaction is a well-known prebiotic pathway for synthesizing ribose from formaldehyde (CH2O) breslow1959mechanism . Our simulations have generated a few relevant hydrocarbon products (shown in Figures S8 and S9), including formaldehyde molecule (number 1 in Figure S8), so we chose it as the primary reactant of synthesizing ribose. Additionally, to facilitate the synthesis process, the reactants also include a “C2” molecule, (Z)-ethene-1,2-diol (C2H4O2, number 7 in Figure S8). Initially, We constructed a simulation box consisting of 10 (Z)-ethene-1,2-diol molecules and 5 formaldehyde molecules (Figure S10). After about 76 ps, a “C5” molecule was formed at $\sim$ 10 GPa and 1400 K. However, it could not transform into the structure of ribose after an extended time. We further inserted 5 water molecules to the system, and finally got ribose molecules. The reaction process is shown in Fig. 5 (a), where formaldehyde and “C2” molecules reacted to form a “C3” molecule, which subsequently combined with another “C2” molecule to form a “C5” molecule. This “C5” molecule then evolved into the open-chain form of ribose at about 19 ps.
To analyze the formation process, we calculated the free energy in Fig. 5 (b), where two C-C distances were selected as collective variables labelled as CV1 and CV2. For the first collective variable, two carbon atoms come from formaldehyde and “C2” molecule, respectively, which are labeled as C1 with yellow color. While for the second collective variable, two carbon atoms are from two different “C2” molecules, marked as C2 with orange color. Our aim is to determine whether ribose formation follows the C1-C1-C2-C2 connectivity pattern. The free energy landscape shows a clear energy minimum that corresponds to the formation of ribose. However, the C-C distance at this point is larger than the typical C-C bond distance, indicating that the formation of ribose deviates from our initial expectations. To gain further insights, we examined the biased molecular dynamics trajectories and found that the carbon backbone of generated ribose molecule follows the C1-C2-C1-C2 arrangement. By measuring the distances of C1-C1 and C2-C2 in this molecule, we confirm that they correspond to the distances observed at the minimum energy point. Furthermore, an alternative initial configuration for the free energy calculation is also presented in Figure S11. Despite the unexpected bondings of the carbon atoms, the free energy landscape remains a reliable indicator of the ribose formation process, as the region of minimum energy still corresponds to a formation of ribose molecule.
Previous studies showed that the cyclic forms of ribose are abundant in aqueous solutions at room temperature drew199813c . However, the cyclic forms of ribose were not detected after 90 ps in our simulations (Fig. 5 (a)). We also carried out further simulations to try to generate rings, as illustrated in Figure S12, but there was still no ring formed. However, that does not mean that the cyclic forms of ribose could not be formed under extreme conditions. The simulation time of 90 ps, though relatively long for AIMD simulations, is significantly shorter than the timescales of real chemical evolutions in deep Earth. The spacial limitation under high pressure may restrict the configuration changes of ribose, reducing the probability of the C4’ or C5’ hydroxyl group attacking the C1’ aldehyde group, which is disadvantageous for producing the furanose form (5-membered ring) or pyranose form (6-membered ring). These two factors suggest that some cyclic formation reactions may be restricted by the shorter timescale and small size of the system in AIMD simulations. To better study the cyclic formation mechanism, we chose the curved open-chain ribose molecules obtained from the previous simulations, as it appears that the C4’ or C5’ hydroxyl group has a greater chance of reacting with C1’ aldehyde group under such narrow space. We built a new simulation box containing three of these ribose molecules and eighteen water molecules. After about 36ps, a furanose form of ribose was formed. Fig. 5 (c) shows the ring formation mechanism, the C4’ hydroxyl group first lost a hydrogen atom, leaving a carbonyl group. The oxygen atom then attacked C1’ to form the ring, and finally, the carbonyl group on C1’ received one dissolved proton in the solution to generate the 5-membered ribose ring. No 6-membered ring was found from our simulations of extreme P-T conditions.
We further applied free energy calculations to study the formation ability of 5- and 6-membered ribose rings at $\sim$ 10 GPa and 1400 K. We chose two C-O distances as the CVs for the two rings, as shown in Fig. 5 (d). The energy minima are at the C-O distance of 1.37 Å for the two ribose forms, indicating the formation of C-O bonds and the rings. Besides, the formation of the C-O bonds lowers the free energy by 11.57 kcal/mol for the 5-membered ribose ring and 5.85 kcal/mol for the 6-membered ribose ring , indicating that the ring formation is thermodynamically more stable and the furanose form of ribose is more favored than the pyranose form at $\sim$ 10 GPa and 1400 K. The free energy result explained why the the pyranose form of ribose is absent in our AIMD simulations. Compared with ambient conditions, high pressure would induce the limited space, which may not support the existence or formation of bigger cyclic form of molecules, so energy barrier for forming 5-membered ring is lower than that of 6-membered ring.
Note that at ambient conditions, the naturally-occurring ribose exists as a mixture of cyclic forms in equilibrium with its open-chain form in aqueous solutions, and the pyranose form of ribose is the predominant conformation, accounting for approximately 70% dewick2013essentials ; dass2021equilibrium . However, despite the thermodynamic stability of the pyranose form of ribose over the furanose form at ambient conditions, the exclusive constituent of the carbohydrate backbone of RNA remains the furanose form of ribose. The reason behind it is still unknown. There are several hypotheses proposed to explain this preference, including silicate(borate)/ribose complexes kolb2004complexes ; lambert2004silicate and temperature gradients dass2021equilibrium . Our finding suggests that extreme P-T conditions in Earth’s upper mantle may also generate the necessary component for assembling RNA molecules. Consequently, it highlights the potential role of upper mantle environment in facilitating the emergence and development of RNA-based life.
We then shifted our attention towards the chemical evolution of nucleobases, which are another essential building blocks of genetic information. There are five naturally-occurring nucleobases: adenine (A), cytosine (C), guanine (G), thymine (T), and uracil (U). Of these, uracil is unique to RNA, so our studies are primarily focused on investigating the formation of uracil molecules under extreme conditions. We first studied the formation of the open-chain “C4N2” molecule with the same CN-backbone present in uracil, as illustrated in Fig. 6 (a). Previous experimental studies on the origins of life have suggested that formamide is a crucial molecule for the prebiotic formation of nucleobases, in the presence of mineral or metal oxide catalysts saladino2012formamide , heatingenchev2021chemical , and high-energy free radicals ferus2015high ; ferus2014high , so here we also first chose formamide as a reactant. However, after about 144 ps, we did not observe the formation of the “C4N2” molecule at 10 GPa and 1400 K, indicating that it is difficult for small molecules to directly synthesize uracil in the AIMD simulations at extreme P-T conditions (Figure S13). Later, we chose the chemical reactants that have closer structural similarities to the “C4N2” molecule. The reactant molecules, including “C4N”, “C3N2”, “C2N2”, and “C3”, were generated from our previous AIMD simulations. One previous study suggested that uracil could be synthesized from the reactions between urea and “C3” molecules choe2022mechanism , so we specifically selected the C3H4O4 molecule (number 10 in Figure S8) and the urea molecule (CH4N2O, number 11 in Figure 1) as reactants. The simulation results are presented in Figure S14. However, regardless of the inclusion of water molecules or the application of different pressure conditions, we do not observe any “C4N2” molecules with the CN-backbone of uracil. We conducted eight more AIMD simulations to investigate the possibility of forming uracil through reactions involving those molecules at various pressures and 1400 K (see the Supporting Information).
[…]
To study the ring formation of the uracil molecule, we also selected the “C4N2-1” molecule with large curvature and built a simulation box containing 6 “C4N2-1” and 6 water molecules. Fig. 6 (b) shows that all the “C4N2-1” molecules reacted. At 5.6 GPa and 1400 K, linear “C4N2-1” molecules became 5-membered and 6-membered rings within a shorter time of about 3.6 ps. Particularly, the elements in the six-membered-ring molecule are the same as those in uracil as shown in Fig. 6(a). At 9.3 GPa and 1400 K, the same 6-membered-ring molecule is also formed. These simulations suggest that it is possible for the open-chain form of molecules with same CN-backbone of uracil to transform into the cyclic form. We further performed free energy calculations to analyze this ring formation reaction. We first chose the distance between the head and tail carbon atoms in the open-chain form of uracil as the CV. The enhanced sampling simulation at $\sim$ 10 GPa and 1400 K produced the same 6-membered-ring molecule as in our previous simulation. Note that the free energy profile suggests that the cyclic form is at the metastable state.
Later we chose allophanic acid (C2H4N2O3, number 13 in Fig. 1 ) and (Z)-ethene-1,2-diol (C2H4O2, number 7 in Figure S8) as reactants, and built a new simulation box containing 5 C2H4N2O3, 5 C2H4O2, and 10 H2, as shown in Fig.6 (c). At $\sim$ 10 GPa and 1400 K, a “C5N2” molecule, 3-(3-carboxyureido)-2,3-dihydroxypropanoic acid, was formed at around 28.8 ps, This molecule can be considered as the CN-backbone of thymine (T). As the simulation continued, a “C4N2” molecule, 2,3-dihydroxy-3-ureidopropanoic acid, labeled as “C4N2-2”, was formed. The formation process of these two molecules are shown in Figure S18. We further performed a enhanced sampling simulation to analyze the ring formation process at $\sim$ 10 GPa and 1400 K. The simulation box has one “C4N2-2” molecule and 32 water molecules, as shown in Fig. 6 (c). Finally, the molecule with a uracil ring, known as uracil glycol, was successfully synthesized, which is a uracil-like molecule. The free energy curve in Fig. 6 (c) shows that the cyclic form is more stable than the open-chain form, indicating that this uracil-like molecule is possible to form in geological C-H-O-N fluids at extreme P-T conditions. we do not postulate that uracil molecules cannot be synthesized due to the limited time scale and length scale of such atomistic simulations. As molecules containing the uracil ring have already been observed, we are confident that with a longer duration of chemical reactions, it is highly probable to produce complete uracil molecules in geological fluids. The uracil molecule can then serve as a precursor for the synthesis of other three types of nucleobases, providing the necessary components for forming RNA molecules.
III Conclusions
Extensive AIMD simulations have been performed to investigate the chemical reactions of small inorganic molecules under the extreme conditions found in Earth’s upper mantle. Three corresponding P-T systems were considered, i.e., 10 GPa - 1000 K, 10 GPa - 1400 K, and 13 GPa - 1400 K. The simulation results have revealed the potential formation of substantial quantities of organic matter, with the composition and abundance of CN-containing species being profoundly influenced by the specific P-T conditions. Notably, formamide, recognized as a crucial precursor in the production of biomolecules, is observed to readily form under high P-T conditions, and its formation pathways are compared across different P-T conditions. Furthermore, free energy calculations have confirmed that the 10 GPa - 1400 K condition is particularly favorable for the generation and preservation of CN-containing compounds.
Using the molecules formed, we further explore the formation process of large biomolecules that are directly related to the building blocks of life. Remarkably, glycine, ribose, and uracil-like molecules are successfully synthesized under extreme conditions in our AIMD simulations, thereby providing initial insights into the formation of organic compounds critical to the emergence of life. More importantly, different from that the 6-membered ring of ribose is the majority in solutions at ambient conditions, it is found that formation of 5-membered ring form of ribose is more favorable at Earth’s mantle conditions, which could provide the necessary components for the generation of RNA molecules. Moreover, these compounds have the potential to be transported to the Earth’s surface through geological activities, such as volcanic eruption, consequently contributing to the origins of life. Our comprehensive studies present a novel perspective suggesting that life may have originated from the depths of the Earth’s interior. The findings underscore the profound impact of extreme P-T conditions on the formation of organic matter and provide valuable insights into the mechanisms underlying the emergence of life on our planet.
Methods
III.1 First Principles molecular dynamics simulations (FPMD)
Born-Oppenheimer ab initio molecular dynamics (AIMD) simulations are conducted using the Qbox codegygi2008architecture to investigate the reactions of C-H-O-N fluids under Earth’s upper mantle conditions. The PBE exchange-correlation functional and ONCV pseudopotentials hamann2013optimized are utilized, and the Bussi-Donadio-Parrinello (BDP) thermostat, with a relaxation time of 24.2 fs, is employed to control the temperature bussi2007canonical . To facilitate a larger time step of 0.24 fs, deuterium atoms are used in place of hydrogen atoms in simulations. For the initial setup, the simulation box consists of 15 NH3, 15 H2O, 15 H2, and 15 CO, as depicted in Figure S1. Periodic boundary condition are applied. Three P-T systems are considered based on Earth’s upper mantle conditions: 10 GPa - 1000 K, 10 GPa - 1400 K, and 13 GPa - 1400 K. NPT simulations are carried out initially to achieve the desired system pressure, with the plane-wave energy cutoff set to 85 Ry for the real cell and 70 Ry for the reference cell. The results obtained from NPT simulations are presented in the supporting information (Figure S2). Once the size of the unit cell is converged, a new box will be built based on the equilibrium stage in NPT simulations to run NVT simulations (Table S1). The energy cutoff is adjusted to 65 Ry. Simulations spanning a duration of over 400 ps are conducted under different P-T conditions to ensure the attainment of system equilibrium. All chemical product analyses presented in the main text are obtained from the NVT simulations. Figure S3 shows the total energy and temperature variations during the simulation period. Within the simulation time, the total energy gradually changes little, and the temperature remains stable in all systems. Pressure is also monitored in NVT simulations. To improve computational efficiency, pressure values are not directly imported from NVT simulations but instead calculated by carrying out independent simulations at different time points of NVT simulations with an energy cutoff of 85 Ry. We select random time points under different P-T conditions for the test, which reach approximately 200 ps of NVT simulations, and every data point of pressure is obtained by averaging 12,000 steps, as shown in Figure S4. The total and component values are all within the error ranges. Therefore, we can confidently assert that our NVT simulations are executed under the target P-T conditions.
Free-energy calculations by coupling FPMD simulations with enhanced sampling
To compute the free-energy landscape for compound formation, we employ the adaptive biasing force method (ABF) darve2008adaptive as implemented in the software package SSAGES sidky2018ssages coupled with the FPMD software Qbox. Different from other advanced sampling methods that add biases to the energy of the system, this method seeks a flattening of the generalized force. The basic ABF algorithm is shown as follows, the running average of the force in bin k along the reaction coordinate space $\xi$ could be calculated:
${{F}_{\xi}}\left({{N}_{step,}}k\right)=\frac{1}{\nu\left({{N}_{step,}}k\right)}\sum\limits_{i=1}^{\nu\left({{N}_{step,}}k\right)}{{{F}_{i}}}\left(t_{i}^{k}\right)$
${{F}_{i}}(t_{i}^{k})=\frac{d}{dt}\left({{M}_{\xi}}\frac{d\xi}{dt}\right)\left|{}_{t_{i}^{k}}\right.$
[…]
In free-energy calculations, similar to the FPMD simulations described above, we utilize enhanced sampling simulations in the NVT ensemble, employing the BDP thermostat with a time step of 0.24 fs, using the PBE exchange-correlation functional with plane waves (a 65 Ry energy cutoff), and the ONCV pseudopotential. For each specific calculation, we use different numbers of grids for distance variables to store local estimates of the mean derivative of the free energy surface. Furthermore, we select different kinds of distance between atoms as collective variables (CVs) for different organic compounds.
Supporting information
Details of simulations, percentage of the C-N containing species with time, hydrocarbon products and free energy calculation, another strategies of formation of ribose and uracil, and their rings.
Acknowledgements
The authors would like to acknowledge support from the Croucher Foundation through the Croucher Innovation Award, Hong Kong Research Grants Council (Projects GRF-16307618, GRF-16306621, and C6021-19EF), National Natural Science Foundation of China (Project 11774072 and Excellent Young Scientists Fund), the Alfred P. Sloan Foundation through the Deep Carbon Observatory, and the Hetao Shenzhen/Hong Kong Innovation and Technology Cooperation (HZQB-KCZYB-2020083).

### v2

@@
-Synthesis and stability of biomolecules under Earth’s upper mantle conditions
+Present address: ]Lehrstuhl für Theoretische Chemie, Ruhr-Universität Bochum, 44780 Bochum, Germany
 Abstract
-How life started on Earth is a long-time unsolved mystery.
-There are various hypotheses ranging from outer space to seabed.
-Here, we applied extensive ab initio molecular dynamics (AIMD) simulations to study chemical reactions of NH3, H2O, H2, and CO at pressures (P) and temperatures (T) approximating the conditions of Earth’s upper mantle (i.e.
+How life started on Earth is an unsolved mystery.
+There are various hypotheses for the location ranging from outer space to the seafloor, subseafloor or potentially deeper.
+Here, we applied extensive ab initio molecular dynamics (AIMD) simulations to study chemical reactions between NH3, H2O, H2, and CO at pressures (P) and temperatures (T) approximating the conditions of Earth’s upper mantle (i.e.
 10 - 13 GPa, 1000 - 1400 K).
-Contrary to the previous assumption that larger organic molecules might readily dissociate in aqueous solutions at extreme P-T conditions, we found that many organic compounds formed and persisted in C-H-O-N fluids under these extreme conditions, including glycine, ribose, and uracil-like molecules.
+Contrary to the previous assumptions that larger organic molecules might readily disintegrate in aqueous solutions at extreme P-T conditions, we found that many organic compounds formed without any catalysts and persisted in C-H-O-N fluids under these extreme conditions, including glycine, ribose, urea, and uracil-like molecules.
 Particularly, our free energy calculations showed that the C-N bond is thermodynamically stable at 10 GPa and 1400 K.
-Moreover, our findings support the “RNA world” hypothesis, as we observed the exclusive formation of the 5-membered-ring form of ribose.
-By exploring the depths of Earth’s interior, we have uncovered a previously unexplored pathway through which life may have originated.
-These findings have contributed to our evolving understanding of the fundamental conditions necessary for life to arise on our planet.
+Moreover, while the pyranose (six-membered-ring) form of ribose is more stable than the furanose (five-membered-ring) form at ambient conditions, we observed the predominant formation of the five-membered-ring form of ribose at extreme conditions, which is consistent with the exclusive incorporation of $\beta$-D-ribofuranose in RNA.
+We have uncovered a previously unexplored pathway through which the building blocks of biomolecules may have originated in early Earth and other planets.
+Our findings contribute to an evolving understanding of the fundamental conditions necessary for life to arise.
 I Introduction
 The origin of life on Earth remains a profound scientific mystery.
-Numerous hypotheses have been put forward, but none have been conclusively verified perez2020prebiotic .
-As early as $\sim$150 years ago, Darwin proposed a “warm little pond” idea darwin1871jd , which was later developed by Oparin and Haldane into the influential ”primordial soup” theory in the 1920s oparin1957origin , that is, the reactions of small molecules, such as CH4, NH3, H2O, and CO2, leads to the abiotic origin of life on early Earth.
+Numerous hypotheses have been put forward, but not a single one can explain all orgel1994origin ; mann2013origins ; sutherland2017opinion .
+As early as $\sim$150 years ago, Darwin proposed a “warm little pond” idea darwin1871jd , which was later developed by Oparin and Haldane into the influential “primordial soup” theory in the 1920s oparin1957origin , that is, the reactions of small molecules, such as CH4, NH3, H2O, and CO2, to form the first organic compounds.
+This leads to the abiotic origin of life on early Earth.
 In 1953, the famous Miller-Urey experiment was conducted to test this hypothesis at the presence of electric discharge miller1953production .
-Inspired by the ”primordial soup” theory, various prebiotic conditions have been proposed for the origin of life, ranging from the interstellar medium (ISM) herbst1995chemistry or other planets burton2012understanding ; sleep1998refugia to deep-sea hydrothermal vents takai2006ultramafics .
-Hydrothermal vents are fissures in Earth’s oceanic crust where hot water rich in gases and minerals is released.
-Those deep sea vents are considered to have suitable conditions for origin of life.
-The pressure (P) of hydrothermal vents is tens of MPa depending on the depth of water, and the temperature (T) can reach 700 K.
-It is widely believed that the origin of life is unlikely to occur in the deeper regions of the Earth, such as the upper mantle, where the pressure can reach up to $\sim$13 GPa, accompanied by temperatures of $\sim$1700 K; the extremely high P-T conditions could readily degrade crucial biomolecules in aqueous solutions miller1988submarine .
-Otake et al.
-studied the stability of amino acids and their oligomerization at 10–5.0 GPa and 453–673 K, and found that high pressure inhibited the decomposition of amino acids ; however, their experiments were conducted at the water-poor condition otake2011stability .
-Previously, the fluids in Earth’s mantle were commonly modelled as simple mixtures of small volatile molecules, including H2O, CO2, CO, CH4, and H2 zhang2009model .
-However, recent experimental and theoretical studies have revealed the significant roles played by chemical speciation, aqueous ions, and complexes in the supercritical geofluids within Earth’s lithosphere manning2013chemistry ; sverjensky2020changing ; pan2016fate ; stolte2019large ; stolte2021water .
-These findings have challenged the previous assumptions and shed light on the potential existence of larger organic molecules.
-It is very challenging to study the origin of life in laboratory settings due to the extreme environments involved.
-Atomistic simulations based on first principles have been applied to study prebiotic chemistry, offering significant molecular insights perez2020prebiotic .
+Inspired by the “primordial soup” theory, various prebiotic environments have been proposed for the origin of life, ranging from the outer space ziurys2006chemistry ; sandford2020prebiotic ; krasnokutski2022pathway to deep-sea hydrothermal vents daniel2006origins ; martin2008hydrothermal .
+In those hypotheses, prebiotic chemistry typically happens at non-ambient conditions.
+The pressure (P) of hydrothermal vents, resulting from the interaction between seafloor and seawater producing hot fluids, is tens of megapascals depending on the depth of water, and the temperature (T) can reach 700 K.
+Those deep sea vents were considered to have suitable conditions for the origin of life martin2008hydrothermal , but previous studies also suggested that the high temperatures in the vents could readily degrade crucial biomolecules in aqueous solutions miller1988submarine .
+Shock recovery experiments successfully synthesized various biomolecules, where the pressure can reach tens of gigapascals and the temperature can rise to thousands of kelvins before releasing back to ambient conditions Sekine2024 ; however, due to the dynamic nature of shock waves, samples do not reach thermodynamic equilibrium states, so the shock experiments are usually conducted to mimic the conditions of meteorite impacts on Earth furukawa2009biomolecule .
+It was commonly assumed that large biomolecules could not persist at both high hydrostatic pressures and high temperatures as found in Earth’s upper mantle, where pressures can reach up to $\sim$13 GPa, accompanied by high temperatures of $\sim$1700 K.
+In recent years, Otake et al.
+studied the stability of amino acids and their oligomerization at 1.0–5.5 GPa and 453–673 K, and found that high-pressure conditions inhibited the decomposition of amino acids and small amino acids could be oligomerized up to pentamers; however, their experiments were conducted at the water-poor condition otake2011stability .
+In aqueous solutions, there seems a strong thermodynamic drive towards the formation of amino acids dimers above 1.5 GPa and temperature above 423 K robinson2021quantifying , in agreement with some experimental observations pedreira2019spontaneous .
+Previously, the fluids in Earth’s mantle were commonly modeled as simple mixtures of small unreactive volatile molecules, including H2O, CO2, CO, CH4, and H2 zhang2009model .
+However, recent experimental and theoretical studies have revealed the significant roles played by chemical speciation, aqueous ions, and complexes in the supercritical geofluids within Earth’s lithosphere manning2013chemistry ; pan2013dielectric ; sverjensky2020changing ; pan2016fate ; abramson2017water ; stolte2019large ; stolte2021water .
+These findings have challenged the previous assumptions and shed light on the potential existence of larger organic molecules in Earth’s or other planets’ subsurface or even deeper.
+It is very challenging to study the origin of life in laboratory settings due to the extreme environments and subsequent analyses.
+Atomistic simulations based on first principles have been applied to study prebiotic chemistry, offering significant molecular-level insights perez2020prebiotic .
 Saitta et al.
@@
 applied AIMD simulations to investigate the formation of prebiotic compounds from impact-induced shock compression of cometary ices followed by expansion to ambient conditions, and found that the impact of cometary ices could produce amino acids independently of atmospheric conditions and Earth minerals goldman2010synthesis ; goldman2013prebiotic ; koziol2015prebiotic .
+Note that the shock compression duration was less than 10 ps, so their simulations did not reach thermal equilibrium, and it is unknown whether the biomolecules can remain stable without being quenched to ambient conditions.
 AIMD simulations coupled with enhanced sampling techniques have also been applied to study the role of mineral surfaces in prebiotic chemistry nair2006glycine ; pollet2006ab ; schreiner2011peptide .
-Despite the numerous previous studies, the exploration of prebiotic chemistry within Earth’s upper mantle remains an uncharted territory.
-Here, we applied extensive AIMD simulations to study the C-H-O-N fluids at 10–13 GPa and 1000–1400 K, the P-T conditions as found at the bottom of Earth’s upper mantle.
-We discovered a large range of organic molecules directly formed without any catalysts at extreme P-T conditions, including glycine, ribose, and uracil-like molecules.
+Despite the numerous previous studies, the exploration of prebiotic chemistry at both high hydrostatic pressures and high temperatures largely remains an uncharted territory.
+Here, we applied extensive AIMD simulations to study the C-H-O-N fluids at 10–13 GPa and 1000–1400 K, the P-T conditions as found in the transition zone of Earth’s upper mantle.
+We discovered that a large range of organic molecules directly formed without any catalysts at extreme P-T conditions, including formamide, glycine, ribose, urea, and uracil-like molecules.
 Our free energy calculation suggested that the C-N bond is thermodynamically stable at 10 GPa and 1400 K.
-Our study has provided valuable insights into a new potential avenue for the origin of life within Earth’s interior, expanding our understanding of the conditions necessary for life to emerge.
+Our study provides valuable insights into a new potential avenue for the formation of building blocks of life and potentially for the origin of life within Earth’s and other planets’ interior, expanding our understanding of the conditions necessary for life to emerge.
 II Results and Discussion
-We initially mixed the H2O, H2, CO, and NH3 molecules in the simulation box.
-The molecular number of each molecule is 15.
-We conducted AIMD simulations at three P-T conditions: 10 GPa and 1000 K, 10 GPa and 1400 K, and 13 GPa and 1400 K to see how chemical products vary with P and T.
+The simulation box initially contained 15 molecules of each of the following species: H2O, H2, CO, and NH3 (Figure S1).
+Those molecules are typical volatile compounds found in C-H-O-N fluids in reducing environments of Earth’s deep interior zhang2009model ; manning2013chemistry , and are also abundant in interstellar space guelin2022organic .
+We conducted AIMD simulations at three P-T conditions, 10 GPa and 1000 K, 10 GPa and 1400 K, and 13 GPa and 1400 K, to evaluate how chemical products vary with P and T.
 Those P-T conditions are typically found at the bottom of Earth’s upper mantle.
-After 400 ps simulations at each P-T condition, we found about 100 different organic species.
-Here we mainly focus on the formation of bio-relevant molecules, so we classify the organic molecules into six categories: “CN”, “C2N”, “CN2”, “C2N2”, “C3Nx”, and “C4Nx”, based on the numbers of carbon and nitrogen atoms, as shown in Table 1.
+After more than 400 ps simulations at each P-T condition, we found about 100 different organic species.
+Here, we mainly focus on the formation of biotically relevant molecules, so we classify the organic molecules into six categories: “CN”, “C2N”, “CN2”, “C2N2”, “C3Nx”, and “C4Nx”, based on the numbers of carbon and nitrogen atoms, as shown in Table 1.
 Fig.
 1 shows the chemical structures of the representative species.
-Species with more than five carbon atoms are excluded due to their high instability and extremely short lifetime, typically lasting for only a few femtoseconds.
-The number and percentage of these CN-containing species as functions of time under three different P-T conditions are shown in Figures S5 and S6, respectively, which suggest that the concentrations of these species gradually stabilize over time.
+Species with more than five carbon atoms were observed; however they are excluded at this stage due to their high instability and extremely short lifetime, typically lasting for only a few femtoseconds.
+The percents of these CN-containing species as a function of time at three different P-T conditions are shown in Figure S3, which suggest that the concentrations of these species gradually stabilize over time.
 At 10 GPa and 1000 K, most of the organic products belong to three main categories, as shown in Fig.
 2(a).
-The “C2N2” species account for more than half of the CN-containing species, followed by the “CN” species, and finally “C2N”.
+The “C2N2” species accounts for more than half of the CN-containing species, followed by the “CN” species, and finally “C2N” species.
 Formamide is the primary product in the “CN” species saladino2012formamide .
 It contains the peptide bond, which is a significant precursor for the production of more complex biomolecules, such as amino acids pietrucci2015formamide , nucleic acids, proteins, and even sugars saladino2012genetics ; halfen2011formation .
-The “C2N” species has the same CN-backbone (“-N-C-C-”) as that of glycine, suggesting that the “C2N” species may be a potential reactant for the formation of amino acids.
-At this P-T condition, some of the “C2N” species are not stable and can be transformed into the “C2N2” species by reacting with ammonia or ammonium ions, so the CN-backbone of “C2N2” is “-N-C-C-N-”, and one interesting molecule was identified as alpha-hydroxy-glycineamide ($\alpha$-HGA).
-With increasing the temperature to 1400 K along the isobar at 10 GPa, all the six categories of organic products at this P-T condition are observed.
-The “CN2” and “C4Nx” fractions are very small, while the other four species have similar fractions.
-Compared with 10 GPa and 1000 K, the portion of “C2N2” becomes smaller, while there are more “C2N” species.
-The “C2N” species formed here also share the same CN backbone (“-N-C-C-”) with glycine.
-Notably, in the “C2N” species, it is possible to form $\alpha$-hydroxyglycine, which can subsequently evolve into glycine through simple chemical reactions saitta2014miller .
-The “C2N2” species have the CN-backbone of“-N-C-C-N-”, including $\alpha$-HGA and oxamide.
+The “C2N” species has the same CN-backbone (“-N-C-C-”) as glycine, suggesting that the “C2N” species may be a potential reactant for the formation of amino acids.
+At this P-T condition, some of the “C2N” species are not stable and can be transformed into the “C2N2” species by reacting with ammonia or ammonium ions, so the CN-backbone of “C2N2” species is “-N-C-C-N-”, and one interesting molecule is identified as alpha-hydroxy-glycineamide ($\alpha$-HGA).
+With increasing the temperature to 1400 K along the isobar at 10 GPa, all the six categories of organic products at this P-T condition were observed.
+Fractions of the “CN2” and “C4Nx” species are very small, while the other four species have similar fractions.
+Compared with 10 GPa and 1000 K, the portion of “C2N2” species becomes smaller, while there are more “C2N” species.
+The “C2N” species formed here also shares the same CN-backbone (“-N-C-C-”) with glycine.
+Notably, in the “C2N” species, we found a molecule of $\alpha$-hydroxyglycine, which can easily evolve into glycine saitta2014miller .
+The “C2N2” species has the CN-backbone of “-N-C-C-N-”, including $\alpha$-HGA and oxamide.
 They are important organic molecules in the formation of the building blocks of peptides ligterink2019formation .
@@
 With increasing the pressure to 13 GPa at 1400 K, the “CN” species becomes dominating, and the “C2N2” species comes next.
-For the “CN” species, in addition to formamide, some other crucial precursors for producing more complex biomolecules are identified, such as carbamic acid tossell2010happens , isocyanic acid belson1982preparation , and formimidic acid pranata1995computational .
-The “C2N2” species exhibit a distinct CN-backbone compared to the two systems at 10 GPa, forming as “-C-N-C-N-”, which is present in the organic matter of allophanic acid.
-This kind of acid plays a vital role in the synthesis of biomolecules (e.g.
-uracil), and would be used in subsequent simulations.
-Surprisingly, the fraction of “C2N” species sharply decreases to a small value, adopting a new CN-backbone structure of “-C-N-C-”, as observed in N-formylformamide sanz2009effects .
+For the “CN” species, besides formamide, some other crucial precursors for producing more complex biomolecules are identified, such as carbamic acid tossell2010happens , isocyanic acid belson1982preparation , and formimidic acid pranata1995computational .
+The “C2N2” species exhibits a distinct CN-backbone compared to the two solutions at 10 GPa, forming as “-C-N-C-N-”, which is present in allophanic acid.
+This acid plays a vital role in the synthesis of biomolecules, e.g., uracil, and would be used as reactants in subsequent simulations of forming larger biomolecules.
+Surprisingly, the fraction of “C2N” species sharply decreases, adopting a new CN-backbone structure of “-C-N-C-”, as observed in N-formylformamide sanz2009effects .
 This structural change makes it challenging for “C2N” species formed here to evolve into glycine-like molecules or larger amino acids.
-In the “CN2” species, we have discovered the presence of urea molecules.
-Urea is closely related with biological processes, so it would also serve as a valuable reactant for investigating the reactions to larger biomolecules.
-For all the three P-T conditions, a common feature is that the “CN” and “C2N2” species are always the major compositions among the six CN-containing species.
+In the “CN2” species, we discovered the presence of urea molecule.
+Urea is closely related with biological processes daniel2006origins and considered to be an important condensation agent in prebiotic chemistry burcar2019stark ; gull2020silicate , so it would also serve as a valuable reactant for investigating the reactions to larger biomolecules.
+For all the three investigated P-T conditions, a common feature is that the “CN” and “C2N2” species are always the major components among the six CN-containing species.
+The lifetime distribution in Fig.
+2 (b) shows that the “CN” and “C2N2” species could exist for a longer time at all three P-T conditions, suggesting that they are relatively more stable than other species.
+Among these generated chemical products, formamide is a crucial precursor molecule that plays a vital role in the synthesis of biomolecules in prebiotic chemistry saladino2012formamide ; pietrucci2015formamide ; enchev2021chemical ; saladino2015meteorite ; rios2015impact ; ferus2015high .
+Thus, we first investigated the formation of formamide molecule.
 Fig.
-2 (b) shows the lifetime of these CN-containing species.
-The lifetime distribution in Fig.
-2 (b) shows that the “CN” and “C2N2” species could exist for a longer time at all the three P-T conditions, suggesting that they are relatively more stable than other species.
-Among these generated chemical products, formamide is a crucial precursor molecule that plays a vital role in the synthesis of biomolecules in prebiotic chemistry saladino2012formamide ; pietrucci2015formamide ; enchev2021chemical ; saladino2015meteorite ; rios2015impact ; ferus2015high .
-Thus, we first investigated the formation of formamide.
-Fig.
-3 (a) provides an illustration of the distinct formation pathways of formamide at three P-T conditions.
-The complete formation processes are shown in Figure S7.
-At 10 GPa and 1000 K, the formation process began with the dissociation of water molecules, generating OH- and H+ ions in the aqueous solution.
-When the H+ ion encountered a CO molecule, a reaction occurred, resulting in the formation of a C-H bond and an aldehyde group (-CHO).
-Subsequently, the aldehyde group reacted with the NH3 molecule, leading to the formation of a C-N bond, and ultimately, one N-H bond broke to produce formamide.
+3 (a) shows the different formation pathways at three P-T conditions, and detailed formation processes are displayed in Figure S4.
+At 10 GPa and 1000 K, the formation reaction begins with the dissociation of water molecules, generating OH- and H+ ions in the aqueous solution.
+The encounter between an H+ ion and a CO molecule leads to the formation of a C-H bond and an aldehyde group (-CH=O).
+Subsequently, the aldehyde group reacts with a NH3 molecule to generate a C-N bond, and then, one N-H bond breaks, ultimately producing formamide.
 At 10 GPa and 1400 K, the formation process differs from that at 10 GPa and 1000 K.
-One N-H bond in NH3 first broke, leading to the formation of a dissociative amino group (-NH2).
-The amino group then reacted directly with a CO molecule, forming the C-N bond.
-Later, the water molecule dissociated again, providing a hydrogen atom that ultimately contributed to the formation of the formamide molecule.
-At 13 GPa and 1400 K, the C-N bond was initially formed through the reaction of NH3 and CO molecules.
-After a very short time, the dissociated hydrogen atom from the water molecule also bonded to the carbon atoms, forming a C-H bond.
-Finally, by releasing a hydrogen atom from the NH3 group, the formamide molecule was generated.
-Although the formation pathways differ at three P-T conditions, we observed two common features: first, the reactants always include CO, H2O, and NH3, and second, the reaction pathways must involve the formation of C-N and C-H bonds and the breakage of N-H bonds.
+One N-H bond in NH3 first breaks, leading to a dissociative amino group (-NH2).
+The amino group then reacts with a CO molecule, forming the C-N bond.
+Later, a water molecule dissociates, providing a proton that ultimately contributes to the formation of the formamide molecule.
+At 13 GPa and 1400 K, the C-N bond is initially formed through the reaction between NH3 and CO molecules.
+After a very short time, a dissociated proton from a water molecule is also bonded to the carbon atom, forming a C-H bond.
+Finally, by releasing a proton from the NH3 group, the formamide molecule is generated.
+Despite the different formation pathways at three P-T conditions, we observed two common features: first, the reactants always have CO, H2O, and NH3, and second, the reactions must involve the formation of C-N and C-H bonds and the breakage of N-H bonds.
 For the generation of biomolecules in living matter, the formation of C-N bonds is the key step.
-We further applied the advanced sampling method (see details in method section) to calculate the free energy profile of the C-N bond formation at three P-T conditions, as shown in Fig.
+We further applied the advanced sampling method (see details in the Methods section) to calculate the free energy profile of the C-N bond formation at three P-T conditions in Fig.
 3(b).
-The collective variable is the distance between the carbon and nitrogen atoms.
-At 10 GPa, 1000 K and 13 GPa, 1400 K, the C-N bonds, whose lengths are 1.33 Å and 1.31 Å  respectively, are at the metastable states.
-The free energy decreases upon dissociation of the C-N bond, indicating that breaking C-N bonds at these two P-T conditions are more favorable thermodynamically.
+The collective variable (CV) is the distance between the carbon and nitrogen atoms.
+Our calculations show that the C-N bonds at 10 GPa, 1000 K and 13 GPa, 1400 K, whose lengths are 1.33 Å and 1.31 Å  respectively, are at the metastable states.
+The free energy decreases upon dissociation of the C-N bond, indicating that breaking C-N bonds at these two P-T conditions are more thermodynamically favored.
 The free energy profiles suggest that the energy barriers of the C-N bond dissociation are 17.57 kcal/mol and 16.40 kcal/mol, at 10 GPa, 1000 K and 13 GPa, 1400 K, respectively.
 As a comparison, at 10 GPa and 1400 K, the free energy increases upon dissociation of the C-N bond (1.31 Å), and the energy barrier is 59.04 kcal/mol, indicating that the C-N bond is thermodynamically and kinetically stable at this P-T condition.
-The free energy result agrees with the data in Fig 2, where there are more CN-containing species at 10 GPa and 1400 K than at the other two P-T conditions.
+The free energy result agrees with the data in Fig.
+2, where there are more CN-containing species at 10 GPa and 1400 K than at the other two P-T conditions.
 Additionally, compared with the free energy profile at 10 GPa and 1000 K, there are two more energy minima at the C-N distance of 2.37 Å  and 2.45 Å at 10 GPa, 1400 K and 13 GPa, 1400 K, respectively.
 The free energy increases upon the dissociation of the C-N pair, indicating that these two minima are thermodynamically stable.
-The C-N distances are larger than that of a typical C-N covalent bond, which strongly implies the presence of an additional atom connecting the two.
+The C-N distances are larger than that of a typical C-N covalent bond, which implies the presence of an additional atom connecting the two.
 This is consistent with a wider variety of organic species at these two P-T conditions.
-Our free energy calculations suggest that the C-N bond formation is favorable at 10 GPa and 1400 K.
-With increasing the pressure to 13 GPa and keeping the temperature at 1400 K, the carbon and nitrogen atoms tend to be bonded together by a third atom.
-After studying the basic chemical products at the three P-T conditions, we now turned to investigate the formation of large biomolecules as potential building blocks of life on Earth.
-Previous simulations have extensively studied glycine, the smallest amino acid, under various conditions saitta2014miller ; goldman2010synthesis ; wang2014discovering .
-In our simulations at 10 GPa and 1400 K, we discovered one molecule very similar to glycine, called $\alpha$-hydroxyglycine.
+Overall, our free energy calculations suggest that the C-N bond formation is favored at 10 GPa and 1400 K.
+After studying the basic chemical products at three P-T conditions, we now turn to investigate the formation of large biomolecules as potential building blocks of life.
+Previous simulations have extensively studied glycine and the smallest amino acid under various conditions saitta2014miller ; goldman2010synthesis ; wang2014discovering .
+In our simulations at 10 GPa and 1400 K, we discovered $\alpha$-hydroxyglycine, very similar to glycine.
 Fig.
 4 (a) illustrates the process of forming $\alpha$-hydroxyglycine from formamide.
-The chemical evolution suggests that, a hydrogen atom attacked the formamide molecule, leading to the creation of a protonated formamide molecule, also known as an aminohydroxymethyl radical.
-Subsequently, a CO2 molecule reacted with the protonated formamide, resulting in the production of a molecule with the “-N-C-C-” CN-backbone through forming a C-C bond.
-Finally, another hydrogen atom was captured by a “-COO” group, ultimately yielding the $\alpha$-hydroxyglycine molecule.
-In comparison to glycine, $\alpha$-hydroxyglycine differs by having one H atom and an “-OH” group instead of two H atoms connecting to the $\alpha$-C.
-Glycine can be obtained if the “-OH” group is replaced by a single H atom.
+A proton attacked the formamide molecule, leading to the creation of a protonated formamide molecule, also known as an aminohydroxymethyl radical.
+Then, a CO2 molecule reacts with the protonated formamide, resulting in the “-N-C-C-” backbone through forming a C-C bond.
+Finally, another proton is captured by a “-COO” group, ultimately yielding the $\alpha$-hydroxyglycine molecule.
+In comparison to glycine, $\alpha$-hydroxyglycine differs by having one hydrogen atom and an “-OH” group instead of two H atoms connecting to $\alpha$-C.
+Glycine can be obtained if the “-OH” group is replaced by a hydrogen atom.
 During this process, the formation of the C-H bond is a crucial step.
-To identify appropriate P-T conditions for further advancing the chemical evolution of glycine, we calculated the free energy as a function of the C-H distance in Fig.
+To explore appropriate P-T conditions for the possible synthesis of glycine, we calculated the free energy as a function of the C-H distance in Fig.
 4 (b).
-Considering that the CN-backbone of the “C2N” species in the 13 GPa - 1400 K is “-C-N-C-”, very different from the CN-backbone in the glycine molecule, we did not consider this system.
-From the energy curves, we can see that the energy minima occurs at 1.07 Å for both 10 GPa - 1000 K and 10 GPa - 1400 K, and the energy barriers for C-H breaking are 21.91 kcal/mol and 22.04 kcal/mol, respectively, which are quite similar.
-However, the change of free energy from C-H bond formation to C-H bond dissociation is more positive at 10 GPa - 1400 K (6.76 kcal/mol) compared with that of 10 GPa - 1000 K (2.63 kcal/mol), which illustrates that 10 GPa - 1400 K condition is still the favorable one for forming C-H bond.
-To test if we can also obtain glycine at extreme P-T conditions, we set up a new simulation box containing 5 $\alpha$-hydroxyglycine, 10 H2, and 10 H2O molecules and perform AIMD simulations at 10 GPa and 1400 K, as shown in Fig.
-4(c).
-Initially, two $\alpha$-hydroxyglycine molecules would connect to a larger molecule, but later transformed into two new molecules after about 28 ps, dehydroglycine and glyoxylic acid.
-We mainly focus on dehydroglycine.
-As the simulation went on, this molecule possessed the amino group when one free proton was bonded to its N atom.
-After a shorter time, the $\alpha$-C also captured one proton, resulting in the final molecular structure of glycine at about 38 ps.
-Additionally, our observations indicate that once glycine is formed at 10 GPa - 1400 K, it remains at a stable state.
+Considering that the CN-backbone of the “C2N” species at 13 GPa - 1400 K is “-C-N-C-”, different from the CN-backbone in the glycine molecule, we did not consider this P-T condition.
+We can see that the free energy minima is at 1.07 Å for both 10 GPa, 1000 K and 10 GPa, 1400 K, and the energy barriers to break the C-H bond are 21.91 kcal/mol and 22.04 kcal/mol, respectively, which are very similar.
+After the C-H bond dissociation, the free energy increases more at 10 GPa and 1400 K (6.76 kcal/mol) than at 10 GPa and 1000 K (2.63 kcal/mol), indicating that the C-H bond is more favored at the former P-T condition.
+To study the reaction from $\alpha$-hydroxyglycine to glycine at extreme P-T conditions, we set up a new simulation box containing 5 $\alpha$-hydroxyglycine, 10 H2, and 10 H2O molecules, as shown in Fig.
+4(c) and performed AIMD simulations at 10 GPa and 1400 K.
+Initially, two $\alpha$-hydroxyglycine molecules combine to form one molecule, but later dissociate to dehydroglycine and glyoxylic acid after $\sim$28 ps.
+As the simulation goes on, the C=N double bond in the dehydroglycine molecule receives two protons and finally dehydroglycine becomes glycine at $\sim$ 38 ps.
 Both proteins and DNA are widely recognized as the important biomolecules for building blocks of life.
 However, the question of which one came first in the emergence of life remains a topic of debate.
-This is due to the fact that the replication of DNA relies on the catalytic activity of enzymes, which are a type of protein.
+This is due to the fact that the replication of DNA relies on the catalytic activity of enzymes, which are a type of proteins.
 Conversely, the production of proteins also needs the guidance of DNA, resulting in a chicken-and-egg dilemma.
-To solve this problem, a famous hypothesis known as “The RNA world” was proposed, which believed that RNA may occur earlier than DNA and protein gilbert1986origin ; nisbet1986origin .
-During this stage, the RNA not only have the ability to reproduce themselves and carry genetic information, but also could catalyze certain biological reactions in life activities like enzymemcclain1987model ; forster1990external .
-So the synthesis of RNA plays an important role in the early stage of the origins of life.
-Generally, RNA is composed of ribose sugar, four kinds of nucleobases, and a phosphate group.
-Considering that our simulation systems only contain C, H, O, and N elements, we then also explore whether the ribose and nucleobases could be synthesized under extreme conditions.
-We first consider Ribose (C5H10O5), composed solely of C, H, O elements.
+To solve this problem, a hypothesis known as “the RNA world” was proposed, which believed that RNA may occur earlier than DNA and protein gilbert1986origin ; nisbet1986origin .
+During this stage, the RNA not only have the ability to reproduce themselves and carry genetic information, but also could catalyze certain biological reactions in life activities like enzymes mcclain1987model ; forster1990external .
+Thus, the synthesis of RNA plays an critical role in the early stage of the origin of life.
+RNA is composed of ribose sugar, four kinds of nucleobases, and a phosphate group.
+Considering that our simulations only involve the carbon, hydrogen, oxygen, and nitrogen elements, we then explored whether the ribose and nucleobases could be synthesized under extreme conditions.
+We first considered ribose (C5H10O5).
 The formose reaction is a well-known prebiotic pathway for synthesizing ribose from formaldehyde (CH2O) breslow1959mechanism .
-Our simulations have generated a few relevant hydrocarbon products (shown in Figures S8 and S9), including formaldehyde molecule (number 1 in Figure S8), so we chose it as the primary reactant of synthesizing ribose.
-Additionally, to facilitate the synthesis process, the reactants also include a “C2” molecule, (Z)-ethene-1,2-diol (C2H4O2, number 7 in Figure S8).
+Our simulations have generated a few relevant hydrocarbon products (shown in Figures S5 and S6), including formaldehyde molecule (number 1 in Figure S5), which was later chosen as the primary reactant.
+Additionally, to facilitate the synthesis process, the reactants also included a “C2” molecule, (Z)-ethene-1,2-diol (C2H4O2, number 7 in Figure S5).
 Initially, we constructed a simulation box consisting of 10 (Z)-ethene-1,2-diol molecules and 5 formaldehyde molecules (Figure S7).
-After about 76 ps, a “C5” molecule was formed at $\sim$ 10 GPa and 1400 K.
-However, it could not transform into the structure of ribose after an extended time.
-We further inserted 5 water molecules to the system, and finally got ribose molecules.
+After $\sim$76 ps, a “C5” molecule was formed at $\sim$10 GPa and 1400 K.
+However, it did not transform into the structure of ribose after an extended time.
+We further tried to add other reactants to the simulation box, and found that after adding 5 water molecules, the ribose molecule appeared, indicating the importance role of water in the formation of ribose.
 The reaction process is shown in Fig.
-5 (a), where formaldehyde and “C2” molecules reacted to form a “C3” molecule, which subsequently combined with another “C2” molecule to form a “C5” molecule.
-This “C5” molecule then evolved into the open-chain form of ribose at about 19 ps.
-To analyze the formation process, we calculated the free energy in Fig.
+5 (a), where the formaldehyde and “C2” molecules react to form a “C3” molecule, which then combines with another “C2” molecule to form a “C5” molecule.
+This “C5” molecule then evolves into the open-chain form of ribose at $\sim$19 ps.
+We calculated the free energy landscape in Fig.
 5 (b), where two C-C distances were selected as collective variables labelled as CV1 and CV2.
-For the first collective variable, two carbon atoms come from formaldehyde and “C2” molecule, respectively, which are labeled as C1 with yellow color.
-While for the second collective variable, two carbon atoms are from two different “C2” molecules, marked as C2 with orange color.
-Our aim is to determine whether ribose formation follows the C1-C1-C2-C2 connectivity pattern.
-The free energy landscape shows a clear energy minimum that corresponds to the formation of ribose.
-However, the C-C distance at this point is larger than the typical C-C bond distance, indicating that the formation of ribose deviates from our initial expectations.
-To gain further insights, we examined the biased molecular dynamics trajectories and found that the carbon backbone of generated ribose molecule follows the C1-C2-C1-C2 arrangement.
-By measuring the distances of C1-C1 and C2-C2 in this molecule, we confirm that they correspond to the distances observed at the minimum energy point.
-Furthermore, an alternative initial configuration for the free energy calculation is also presented in Figure S11.
-Despite the unexpected bondings of the carbon atoms, the free energy landscape remains a reliable indicator of the ribose formation process, as the region of minimum energy still corresponds to a formation of ribose molecule.
+For CV1, two carbon atoms come from formaldehyde and a “C2” molecule, respectively, which are labeled as C1.
+While for CV2, two carbon atoms are from two different “C2” molecules, marked as C2.
+The free energy landscape shows a clear energy minimum corresponding to the formation of ribose, but the C-C distance at this minimum is larger than a typical C-C bond distance.
+We examined the biased molecular dynamics trajectories and found that the carbon backbone of the generated ribose molecule has the C1-C2-C1-C2 structure.
+By measuring the distances of C1-C1 and C2-C2 , we confirmed that they are consistent to the distances observed in the free energy landscape.
+We also randomly selected another initial configuration to confirm that the ribose molecule is thermodynamically stable, as shown in Figure S8.
 Previous studies showed that the cyclic forms of ribose are abundant in aqueous solutions at room temperature drew199813c .
-However, the cyclic forms of ribose were not detected after 90 ps in our simulations (Fig.
+However, the cyclic forms of ribose were not detected in our simulations (Fig.
 5 (a)).
-We also carried out further simulations to try to generate rings, as illustrated in Figure S12, but there was still no ring formed.
-However, that does not mean that the cyclic forms of ribose could not be formed under extreme conditions.
-The simulation time of 90 ps, though relatively long for AIMD simulations, is significantly shorter than the timescales of real chemical evolutions in deep Earth.
-The spacial limitation under high pressure may restrict the configuration changes of ribose, reducing the probability of the C4’ or C5’ hydroxyl group attacking the C1’ aldehyde group, which is disadvantageous for producing the furanose form (5-membered ring) or pyranose form (6-membered ring).
-These two factors suggest that some cyclic formation reactions may be restricted by the shorter timescale and small size of the system in AIMD simulations.
-To better study the cyclic formation mechanism, we chose the curved open-chain ribose molecules obtained from the previous simulations, as it appears that the C4’ or C5’ hydroxyl group has a greater chance of reacting with C1’ aldehyde group under such narrow space.
-We built a new simulation box containing three of these ribose molecules and eighteen water molecules.
-After about 36ps, a furanose form of ribose was formed.
+The limited size of our simulation boxes may restrict the configuration changes of ribose, thereby decreasing the likelihood of the C4’ or C5’ hydroxyl group attacking the C1’ aldehyde group.
+This limitation is unfavorable for the production of the furanose form (five-membered ring) or pyranose form (six-membered ring).
+To further study the cyclic formation mechanism, we carried a few additional simulations and finally chose the curved open-chain ribose molecules obtained from our previous simulations, as it appears that the C4’ or C5’ hydroxyl group has a greater chance of reacting with the C1’ aldehyde group under pressure.
+We built a new simulation box containing three of these ribose molecules together with 18 water molecules.
+After about 36 ps, we found the direct formation of a furanose form of ribose.
 Fig.
-5 (c) shows the ring formation mechanism, the C4’ hydroxyl group first lost a hydrogen atom, leaving a carbonyl group.
-The oxygen atom then attacked C1’ to form the ring, and finally, the carbonyl group on C1’ received one dissolved proton in the solution to generate the 5-membered ribose ring.
-No 6-membered ring was found from our simulations of extreme P-T conditions.
-We further applied free energy calculations to study the formation ability of 5- and 6-membered ribose rings at $\sim$ 10 GPa and 1400 K.
+5 (c) shows the mechanism of ring formation.
+In this process, the C4’ hydroxyl group first loses a hydrogen atom, leaving a carbonyl group.
+The oxygen atom then attacks C1’ to form the ring, and finally, the carbonyl group on C1’ receives a dissolved proton in the solution to generate the five-membered ribose ring.
+Interestingly, no six-membered ring was found in our simulations at 10 GPa and 1400 K.
+We further applied free energy calculations to study the formation ability of five- and six-membered ribose rings at $\sim$ 10 GPa and 1400 K.
 We chose two C-O distances as the CVs for the two rings, as shown in Fig.
@@
 The energy minima are at the C-O distance of 1.37 Å for the two ribose forms, indicating the formation of C-O bonds and the rings.
-Besides, the formation of the C-O bonds lowers the free energy by 11.57 kcal/mol for the 5-membered ribose ring and 5.85 kcal/mol for the 6-membered ribose ring , indicating that the ring formation is thermodynamically more stable and the furanose form of ribose is more favored than the pyranose form at $\sim$ 10 GPa and 1400 K.
+Besides, the formation of the C-O bonds lowers the free energy by 11.57 kcal/mol for the five-membered ribose ring and 5.85 kcal/mol for the six-membered ribose ring, indicating that the ring formation is thermodynamically more stable and the furanose form of ribose is more favored than the pyranose form at $\sim$ 10 GPa and 1400 K.
 The free energy result explained why the the pyranose form of ribose is absent in our AIMD simulations.
-Compared with ambient conditions, high pressure would induce the limited space, which may not support the existence or formation of bigger cyclic form of molecules, so energy barrier for forming 5-membered ring is lower than that of 6-membered ring.
+The spacial limitation under high pressure may hinder the formation of larger cyclic forms of molecules, so the energy barrier of forming a five-membered ring is lower than that of a six-membered ring.
 Note that at ambient conditions, the naturally-occurring ribose exists as a mixture of cyclic forms in equilibrium with its open-chain form in aqueous solutions, and the pyranose form of ribose is the predominant conformation, accounting for approximately 70% dewick2013essentials ; dass2021equilibrium .
@@
 There are several hypotheses proposed to explain this preference, including silicate(borate)/ribose complexes kolb2004complexes ; lambert2004silicate and temperature gradients dass2021equilibrium .
-Our finding suggests that extreme P-T conditions in Earth’s upper mantle may also generate the necessary component for assembling RNA molecules.
+Our finding suggests that extreme P-T conditions may also generate the necessary component for assembling RNA molecules.
 Consequently, it highlights the potential role of upper mantle environment in facilitating the emergence and development of RNA-based life.
-We then shifted our attention towards the chemical evolution of nucleobases, which are another essential building blocks of genetic information.
+We then shifted our attention to the chemical evolution of nucleobases, which are another essential building blocks of genetic information.
 There are five naturally-occurring nucleobases: adenine (A), cytosine (C), guanine (G), thymine (T), and uracil (U).
-Of these, uracil is unique to RNA, so our studies are primarily focused on investigating the formation of uracil molecules under extreme conditions.
-We first studied the formation of the open-chain “C4N2” molecule with the same CN-backbone present in uracil, as illustrated in Fig.
+Of these, uracil is unique to RNA, so our studies primarily focused on investigating the formation of uracil molecules under extreme conditions.
+We first studied the formation of the open-chain “C4N2” molecule which has the same CN-backbone as the open-chain form of uracil, as shown in Fig.
 6 (a).
 Previous experimental studies on the origins of life have suggested that formamide is a crucial molecule for the prebiotic formation of nucleobases, in the presence of mineral or metal oxide catalysts saladino2012formamide , heatingenchev2021chemical , and high-energy free radicals ferus2015high ; ferus2014high , so here we also first chose formamide as a reactant.
-However, after about 144 ps, we did not observe the formation of the “C4N2” molecule at 10 GPa and 1400 K, indicating that it is difficult for small molecules to directly synthesize uracil in the AIMD simulations at extreme P-T conditions (Figure S13).
+However, after about 144 ps, we did not observe the formation of the “C4N2” molecule at 10 GPa and 1400 K, suggesting that it may be difficult for small molecules to directly synthesize uracil in the AIMD simulations (Figure S9).
 Later, we chose the chemical reactants that have closer structural similarities to the “C4N2” molecule.
-The reactant molecules, including “C4N”, “C3N2”, “C2N2”, and “C3”, were generated from our previous AIMD simulations.
-One previous study suggested that uracil could be synthesized from the reactions between urea and “C3” molecules choe2022mechanism , so we specifically selected the C3H4O4 molecule (number 10 in Figure S8) and the urea molecule (CH4N2O, number 11 in Figure 1) as reactants.
+The reactant molecules, including “C4N”, “C3N2”, “C2N2”, and “C3”, were generated from our first-stage AIMD simulations.
+A recent study suggested that uracil could be synthesized from the reactions between urea and “C3” molecules choe2022mechanism , so we specifically selected the C3H4O4 molecule (number 10 in Figure S5) and the urea molecule (CH4N2O, number 11 in Fig.
+1) as reactants.
 The simulation results are presented in Figure S10.
-However, regardless of the inclusion of water molecules or the application of different pressure conditions, we do not observe any “C4N2” molecules with the CN-backbone of uracil.
-We conducted eight more AIMD simulations to investigate the possibility of forming uracil through reactions involving those molecules at various pressures and 1400 K (see the Supporting Information).
+However, regardless of the inclusion of water molecules or the application of different pressure conditions, we did not observe any “C4N2” molecules with the CN-backbone of the open-chain form of uracil.
+Then we tried to find another pathways and conducted six more AIMD simulations to investigate the possibility of forming uracil through reactions involving above-mentioned molecules at various pressures and 1400 K (see Figures S11, S12 and S14).
 […]
-To study the ring formation of the uracil molecule, we also selected the “C4N2-1” molecule with large curvature and built a simulation box containing 6 “C4N2-1” and 6 water molecules.
+To study the ring formation of uracil, we also selected the “C4N2-1” molecule with large curvature and built a simulation box containing 6 “C4N2-1” and 6 water molecules.
 Fig.
 6 (b) shows that all the “C4N2-1” molecules reacted.
-At 5.6 GPa and 1400 K, linear “C4N2-1” molecules became 5-membered and 6-membered rings within a shorter time of about 3.6 ps.
+At 5.6 GPa and 1400 K, open-chain “C4N2-1” molecules became five-membered and six-membered rings within a shorter time of about 3.6 ps.
 Particularly, the elements in the six-membered-ring molecule are the same as those in uracil as shown in Fig.
 6(a).
-At 9.3 GPa and 1400 K, the same 6-membered-ring molecule is also formed.
-These simulations suggest that it is possible for the open-chain form of molecules with same CN-backbone of uracil to transform into the cyclic form.
+At 9.3 GPa and 1400 K, the same six-membered-ring molecule was also formed.
+These simulations suggest that it is possible for the open-chain form of uracil to transform into the cyclic form.
 We further performed free energy calculations to analyze this ring formation reaction.
 We first chose the distance between the head and tail carbon atoms in the open-chain form of uracil as the CV.
-The enhanced sampling simulation at $\sim$ 10 GPa and 1400 K produced the same 6-membered-ring molecule as in our previous simulation.
-Note that the free energy profile suggests that the cyclic form is at the metastable state.
+The enhanced sampling simulation at $\sim$ 10 GPa and 1400 K produced the same six-membered-ring molecule as in our unbiased simulations.
+Note that the free energy profile suggests that the cyclic form is in a metastable state, as shown in Fig.
+6(b).
 Later we chose allophanic acid (C2H4N2O3, number 13 in Fig.
 1 ) and (Z)-ethene-1,2-diol (C2H4O2, number 7 in Figure S5) as reactants, and built a new simulation box containing 5 C2H4N2O3, 5 C2H4O2, and 10 H2, as shown in Fig.6 (c).
-At $\sim$ 10 GPa and 1400 K, a “C5N2” molecule, 3-(3-carboxyureido)-2,3-dihydroxypropanoic acid, was formed at around 28.8 ps, This molecule can be considered as the CN-backbone of thymine (T).
-As the simulation continued, a “C4N2” molecule, 2,3-dihydroxy-3-ureidopropanoic acid, labeled as “C4N2-2”, was formed.
-The formation process of these two molecules are shown in Figure S18.
+At $\sim$ 10 GPa and 1400 K, a “C5N2” molecule, 3-(3-carboxyureido)-2,3-dihydroxypropanoic acid, was formed at around 28.8 ps, This molecule has the same CN-backbone as the open-chain form of thymine (T).
+As the simulation continued, a “C4N2” molecule, 2,3-dihydroxy-3-ureidopropanoic acid, labeled as “C4N2-2”, was formed, which has the same CN-backbone as the open-chain form of uracil.
+The formation process of these two kinds of molecules are shown in Figure S15.
 We further performed a enhanced sampling simulation to analyze the ring formation process at $\sim$ 10 GPa and 1400 K.
@@
 6 (c).
-Finally, the molecule with a uracil ring, known as uracil glycol, was successfully synthesized, which is a uracil-like molecule.
+Finally, a uracil-like molecule, called uracil glycol, was synthesized, which has a uracil ring.
 The free energy curve in Fig.
-6 (c) shows that the cyclic form is more stable than the open-chain form, indicating that this uracil-like molecule is possible to form in geological C-H-O-N fluids at extreme P-T conditions.
-we do not postulate that uracil molecules cannot be synthesized due to the limited time scale and length scale of such atomistic simulations.
-As molecules containing the uracil ring have already been observed, we are confident that with a longer duration of chemical reactions, it is highly probable to produce complete uracil molecules in geological fluids.
-The uracil molecule can then serve as a precursor for the synthesis of other three types of nucleobases, providing the necessary components for forming RNA molecules.
+6 (c) shows that the cyclic form is more stable than the open-chain form, indicating that the uracil-like molecule is possible to form in geological C-H-O-N fluids at extreme P-T conditions.
+The uracil-like molecule can then serve as a precursor for the synthesis of other three types of nucleobases, providing the necessary components for forming RNA molecules.
+Geochemical evidence of biotic activity suggests that the emergence of life can be traced back to as early as 3.8 billion years ago mojzsis1996evidence .
+The modern style of plate tectonics may have started about 3.2 billion years ago brenner2020paleomagnetic , but some form of plate tectonics may have already been active around 4.0 billion years ago.
+and these tectonic activities might carry water, carbon, and other volatile substances down into the mantle korenaga2013initiation ; HOLDER2024 ; blichert2008hafnium ; hopkins2008low ; huang2022barium .
+Because of reducing environments in early Earth’s mantle aulbach2016evidence , oxidized carbon and nitrogen can be reduced to CO and NH3/NH+${}_{4}^{+}$, respectively holm2009reduction (see Figure S16), and they may further synthesize to become key building blocks of life, which may reach the surface of early Earth.
+Our findings also have important implications for organic molecules in interstellar space.
+The reactant molecules H2O, H2, CO and NH3 largely exist in interstellar clouds and circumstellar shells, together with more than 200 complex organic species, such as aldehydes, alcohols, acids, amines and carboxamides guelin2022organic .
+It is unclear how those complex organic species are generated.
+Our current study suggests that the extreme pressure and temperature inside giant planets and their moons neri2020carbonaceous ; nettelmann2016uranus , like Jupiter and Saturn, may help to synthesize those complex organic molecules postberg2018macromolecular ; waite2009liquid ; tosi2024salts .
 III Conclusions
-Extensive AIMD simulations have been performed to investigate the chemical reactions of small inorganic molecules under the extreme conditions found in Earth’s upper mantle.
+Extensive AIMD simulations have been performed to investigate the chemical reactions of small proto-organic molecules under the extreme P-T conditions.
 Three corresponding P-T systems were considered, i.e., 10 GPa - 1000 K, 10 GPa - 1400 K, and 13 GPa - 1400 K.
-The simulation results have revealed the potential formation of substantial quantities of organic matter, with the composition and abundance of CN-containing species being profoundly influenced by the specific P-T conditions.
-Notably, formamide, recognized as a crucial precursor in the production of biomolecules, is observed to readily form under high P-T conditions, and its formation pathways are compared across different P-T conditions.
+The simulation results have revealed the potential formation of substantial quantities of organic matter without any catalysts, with the composition and abundance of CN-containing species being profoundly influenced by the specific P-T conditions.
+Notably, formamide, recognized as a crucial precursor in the production of biomolecules, is observed to readily form under extreme P-T conditions, and its formation pathways are compared across different P-T conditions.
 Furthermore, free energy calculations have confirmed that the 10 GPa - 1400 K condition is particularly favorable for the generation and preservation of CN-containing compounds.
-Using the molecules formed, we further explore the formation process of large biomolecules that are directly related to the building blocks of life.
-Remarkably, glycine, ribose, and uracil-like molecules are successfully synthesized under extreme conditions in our AIMD simulations, thereby providing initial insights into the formation of organic compounds critical to the emergence of life.
-More importantly, different from that the 6-membered ring of ribose is the majority in solutions at ambient conditions, it is found that formation of 5-membered ring form of ribose is more favorable at Earth’s mantle conditions, which could provide the necessary components for the generation of RNA molecules.
-Moreover, these compounds have the potential to be transported to the Earth’s surface through geological activities, such as volcanic eruption, consequently contributing to the origins of life.
-Our comprehensive studies present a novel perspective suggesting that life may have originated from the depths of the Earth’s interior.
-The findings underscore the profound impact of extreme P-T conditions on the formation of organic matter and provide valuable insights into the mechanisms underlying the emergence of life on our planet.
+Using the molecules formed, we have further explored the formation process of large biomolecules that are directly related to the building blocks of life.
+Remarkably, glycine, ribose, urea, and uracil-like molecules are successfully synthesized under extreme conditions in our AIMD simulations, thereby providing initial insights into the formation of organic compounds critical to the emergence of life.
+More importantly, different from that the six-membered ring of ribose is the majority in solutions at ambient conditions, it is found that formation of five-membered ring form of ribose is more favored at extreme P-T conditions, which could provide the necessary components for the generation of RNA molecules.
+Moreover, these compounds have the potential to be transported to Earth’s surface through geological activities, such as mantle convection, consequently contributing to the origins of life.
+The complex organic species found in interstellar space may be also related to the chemical processes inside giant planets and their moons at extreme P-T conditions.
+Our comprehensive studies present a novel perspective extending the cradle of life planetary interiors.
+The findings underscore the profound impact of extreme P-T conditions on the formation of organic matter and provide valuable insights into the mechanisms underlying the emergence of life within planetary bodies.
 Methods
-III.1 First Principles molecular dynamics simulations (FPMD)
-Born-Oppenheimer ab initio molecular dynamics (AIMD) simulations are conducted using the Qbox codegygi2008architecture to investigate the reactions of C-H-O-N fluids under Earth’s upper mantle conditions.
-The PBE exchange-correlation functional and ONCV pseudopotentials hamann2013optimized are utilized, and the Bussi-Donadio-Parrinello (BDP) thermostat, with a relaxation time of 24.2 fs, is employed to control the temperature bussi2007canonical .
-To facilitate a larger time step of 0.24 fs, deuterium atoms are used in place of hydrogen atoms in simulations.
-For the initial setup, the simulation box consists of 15 NH3, 15 H2O, 15 H2, and 15 CO, as depicted in Figure S1.
-Periodic boundary condition are applied.
-Three P-T systems are considered based on Earth’s upper mantle conditions: 10 GPa - 1000 K, 10 GPa - 1400 K, and 13 GPa - 1400 K.
-NPT simulations are carried out initially to achieve the desired system pressure, with the plane-wave energy cutoff set to 85 Ry for the real cell and 70 Ry for the reference cell.
-The results obtained from NPT simulations are presented in the supporting information (Figure S2).
-Once the size of the unit cell is converged, a new box will be built based on the equilibrium stage in NPT simulations to run NVT simulations (Table S1).
-The energy cutoff is adjusted to 65 Ry.
-Simulations spanning a duration of over 400 ps are conducted under different P-T conditions to ensure the attainment of system equilibrium.
-All chemical product analyses presented in the main text are obtained from the NVT simulations.
-Figure S3 shows the total energy and temperature variations during the simulation period.
-Within the simulation time, the total energy gradually changes little, and the temperature remains stable in all systems.
-Pressure is also monitored in NVT simulations.
-To improve computational efficiency, pressure values are not directly imported from NVT simulations but instead calculated by carrying out independent simulations at different time points of NVT simulations with an energy cutoff of 85 Ry.
-We select random time points under different P-T conditions for the test, which reach approximately 200 ps of NVT simulations, and every data point of pressure is obtained by averaging 12,000 steps, as shown in Figure S4.
-The total and component values are all within the error ranges.
-Therefore, we can confidently assert that our NVT simulations are executed under the target P-T conditions.
-Free-energy calculations by coupling FPMD simulations with enhanced sampling
-To compute the free-energy landscape for compound formation, we employ the adaptive biasing force method (ABF) darve2008adaptive as implemented in the software package SSAGES sidky2018ssages coupled with the FPMD software Qbox.
-Different from other advanced sampling methods that add biases to the energy of the system, this method seeks a flattening of the generalized force.
-The basic ABF algorithm is shown as follows, the running average of the force in bin k along the reaction coordinate space $\xi$ could be calculated:
-${{F}_{\xi}}\left({{N}_{step,}}k\right)=\frac{1}{\nu\left({{N}_{step,}}k\right)}\sum\limits_{i=1}^{\nu\left({{N}_{step,}}k\right)}{{{F}_{i}}}\left(t_{i}^{k}\right)$
-${{F}_{i}}(t_{i}^{k})=\frac{d}{dt}\left({{M}_{\xi}}\frac{d\xi}{dt}\right)\left|{}_{t_{i}^{k}}\right.$
+III.1 Ab initio molecular dynamics simulations (AIMD)
+We carried out AIMD simulations with the Born-Oppenheimer approximation using the Qbox code gygi2008architecture .
+We applied the PBE exchange-correlation functional in density functional theory perdew1996generalized .
+It is known that the PBE functional may be insufficient to describe aqueous solutions at ambient conditions, but our previous studies found that it works better at extreme P-T conditions than at ambient conditions pan2013dielectric ; pan2016fate ; stolte2019large ; stolte2021water , and many other studies also showed that this semi-local functional is reliable to study matter under extreme conditions spanu2011stability ; cheng2023thermodynamics ; zhang2023pbe .
+We used the ONCV pseudopotentials hamann2013optimized , and the Bussi-Donadio-Parrinello (BDP) thermostat, with a relaxation time of 24.2 fs, to control the temperature bussi2007canonical .
+To facilitate a larger time step of 0.24 fs, deuterium atoms were used in place of hydrogen atoms in our simulations.
+The simulation boxes have periodic boundary conditions.
+We first performed NPT simulations to achieve the desired pressures, configuring the plane-wave energy cutoff to 85 Ry and the energy cutoff for stress confinement potential to 70 Ry.
+We applied the simulation box sizes (Table S1) obtained from NPT simulations in the following NVT simulations for production runs, where the plane-wave energy cutoff was 65 Ry.
+During NVT simulations, we selected random time points to carry out independent simulations with energy cutoff of 85 Ry to monitor pressure (Figure S2).
+Our total AIMD simulation time exceeds 2.5 ns.
+Free-energy calculations by coupling AIMD simulations with enhanced sampling
+To compute the free-energy landscape, we employed the adaptive biasing force method (ABF) darve2008adaptive as implemented in the software package SSAGES sidky2018ssages coupled with Qbox code.
+Unlike other advanced sampling methods that introduce biases into the energy, this method aims to flatten the generalized force.
+The running average of the force in the $k$th bin along the reaction coordinate $\xi$ is calculated as:
+$F_{\xi}(N_{\text{step}},k)$ $=\frac{1}{\nu(N_{\text{step}},k)}\sum_{i=1}^{\nu(N_{\text{step}},k)}F_{i}(t_{i}^{k})$ (1)
+$F_{i}(t_{i}^{k})$ $=\frac{d}{dt}\left(M_{\xi}\frac{d\xi}{dt}\right)\bigg{|}_{t_{i}^{k}}$ (2)
+where $\nu\left({{N}_{step}},k\right)$ is the number of samples collected in the $k$th bin after $N_{step}$ steps, ${{F}_{i}}\left(t_{i}^{k}\right)$ is the $i$th force sample at the time $t_{i}^{k}$, M is the mass matrix.
+The bias force -$\text{-}\left(\nabla\xi\right){{F}_{\xi}}\left({{N}_{step,}}k\right)$ is added to the calculation until the free energy landscape is flat.
+We finally integrated the total biased force to get the free energy.
+For example, the free energy $\Delta{{A}_{a\to b}}$ between two states $a$ and $b$ is calculated as:
 […]
-[…]
-In free-energy calculations, similar to the FPMD simulations described above, we utilize enhanced sampling simulations in the NVT ensemble, employing the BDP thermostat with a time step of 0.24 fs, using the PBE exchange-correlation functional with plane waves (a 65 Ry energy cutoff), and the ONCV pseudopotential.
-For each specific calculation, we use different numbers of grids for distance variables to store local estimates of the mean derivative of the free energy surface.
-Furthermore, we select different kinds of distance between atoms as collective variables (CVs) for different organic compounds.
 Supporting information
-Details of simulations, percentage of the C-N containing species with time, hydrocarbon products and free energy calculation, another strategies of formation of ribose and uracil, and their rings.
+Details of the simulations, the percentage of C-N containing species over time, the hydrocarbon products, the free energy calculations, other strategies for the formation of ribose and uracil, and their ring structures, compositions of fluids at different oxygen fugacity.
 Acknowledgements
-The authors would like to acknowledge support from the Croucher Foundation through the Croucher Innovation Award, Hong Kong Research Grants Council (Projects GRF-16307618, GRF-16306621, and C6021-19EF), National Natural Science Foundation of China (Project 11774072 and Excellent Young Scientists Fund), the Alfred P.
-Sloan Foundation through the Deep Carbon Observatory, and the Hetao Shenzhen/Hong Kong Innovation and Technology Cooperation (HZQB-KCZYB-2020083).
+This work was supported by the Croucher Foundation through the Croucher Innovation Award, Hong Kong Research Grants Council (Projects GRF-16301723, GRF-16306621, and C6021-19EF), National Natural Science Foundation of China through the Excellent Young Scientists Fund (22022310), and the Hetao Shenzhen/Hong Kong Innovation and Technology Cooperation (HZQB-KCZYB-2020083).
+R.T.
+would like to acknowledge support from the Original Exploration Plan-Key project (42150104).
+Contributions by D.A.S.
+constitute material based upon work supported by the U.S.
+Department of Energy, Office of Science, Basic Energy Sciences, Geosciences program under Award Number DE-SC0019830 as well as NSF Petrology and Geochemistry Grant Number 2032039.
+Part of this work was carried out using computational resources from the National Supercomputer Center in Guangzhou, China.

### v3

@@
 The origin of life on Earth remains a profound scientific mystery.
-Numerous hypotheses have been put forward, but not a single one can explain all orgel1994origin ; mann2013origins ; sutherland2017opinion .
+Numerous hypotheses have been put forward, but not a single one can explain all orgel1994origin ; mann2013origins .
 As early as $\sim$150 years ago, Darwin proposed a “warm little pond” idea darwin1871jd , which was later developed by Oparin and Haldane into the influential “primordial soup” theory in the 1920s oparin1957origin , that is, the reactions of small molecules, such as CH4, NH3, H2O, and CO2, to form the first organic compounds.
@@
 applied AIMD simulations to investigate the formation of prebiotic compounds from impact-induced shock compression of cometary ices followed by expansion to ambient conditions, and found that the impact of cometary ices could produce amino acids independently of atmospheric conditions and Earth minerals goldman2010synthesis ; goldman2013prebiotic ; koziol2015prebiotic .
-Note that the shock compression duration was less than 10 ps, so their simulations did not reach thermal equilibrium, and it is unknown whether the biomolecules can remain stable without being quenched to ambient conditions.
+Note that the shock compression simulations did not reach thermal equilibrium, and it is unknown whether the biomolecules can remain stable without being quenched to ambient conditions Ross2013comment ; Goldman2013reply .
 AIMD simulations coupled with enhanced sampling techniques have also been applied to study the role of mineral surfaces in prebiotic chemistry nair2006glycine ; pollet2006ab ; schreiner2011peptide .
@@
 2 (b) shows that the “CN” and “C2N2” species could exist for a longer time at all three P-T conditions, suggesting that they are relatively more stable than other species.
-Among these generated chemical products, formamide is a crucial precursor molecule that plays a vital role in the synthesis of biomolecules in prebiotic chemistry saladino2012formamide ; pietrucci2015formamide ; enchev2021chemical ; saladino2015meteorite ; rios2015impact ; ferus2015high .
+Among these generated chemical products, formamide is a crucial precursor molecule that plays a vital role in the synthesis of biomolecules in prebiotic chemistry saladino2012formamide ; pietrucci2015formamide ; saladino2015meteorite ; rios2015impact ; ferus2015high ; enchev2021chemical .
 Thus, we first investigated the formation of formamide molecule.
@@
 Previous studies showed that the cyclic forms of ribose are abundant in aqueous solutions at room temperature drew199813c .
-However, the cyclic forms of ribose were not detected in our simulations (Fig.
+However, the cyclic forms of ribose, either furanose form (five-membered ring) or pyranose form (six-membered ring), were not detected in our simulations (Fig.
 5 (a)).
-The limited size of our simulation boxes may restrict the configuration changes of ribose, thereby decreasing the likelihood of the C4’ or C5’ hydroxyl group attacking the C1’ aldehyde group.
-This limitation is unfavorable for the production of the furanose form (five-membered ring) or pyranose form (six-membered ring).
-To further study the cyclic formation mechanism, we carried a few additional simulations and finally chose the curved open-chain ribose molecules obtained from our previous simulations, as it appears that the C4’ or C5’ hydroxyl group has a greater chance of reacting with the C1’ aldehyde group under pressure.
+The configuration change of ribose may require very long simulation time.
+To further study the cyclic formation mechanism, we tried a few additional simulations with new initial states, and finally chose the curved open-chain ribose molecules obtained from our previous simulations, as it appears that the C4’ or C5’ hydroxyl group has a greater chance of reacting with the C1’ aldehyde group under pressure.
 We built a new simulation box containing three of these ribose molecules together with 18 water molecules.
-After about 36 ps, we found the direct formation of a furanose form of ribose.
+After about 36 ps, we observed the direct formation of a furanose form of ribose.
 Fig.
@@
 Our finding suggests that extreme P-T conditions may also generate the necessary component for assembling RNA molecules.
-Consequently, it highlights the potential role of upper mantle environment in facilitating the emergence and development of RNA-based life.
+Consequently, it highlights the potential role of deep Earth environment in facilitating the emergence and development of RNA-based life.
 We then shifted our attention to the chemical evolution of nucleobases, which are another essential building blocks of genetic information.
@@
 6 (a).
-Previous experimental studies on the origins of life have suggested that formamide is a crucial molecule for the prebiotic formation of nucleobases, in the presence of mineral or metal oxide catalysts saladino2012formamide , heatingenchev2021chemical , and high-energy free radicals ferus2015high ; ferus2014high , so here we also first chose formamide as a reactant.
+Previous experimental studies on the origin of life have suggested that formamide is a crucial molecule for the prebiotic formation of nucleobases, in the presence of mineral or metal oxide catalysts saladino2012formamide , heatingenchev2021chemical , and high-energy free radicals ferus2015high ; ferus2014high , so here we also first chose formamide as a reactant.
 However, after about 144 ps, we did not observe the formation of the “C4N2” molecule at 10 GPa and 1400 K, suggesting that it may be difficult for small molecules to directly synthesize uracil in the AIMD simulations (Figure S9).
@@
 The formation process of these two kinds of molecules are shown in Figure S15.
-We further performed a enhanced sampling simulation to analyze the ring formation process at $\sim$ 10 GPa and 1400 K.
+We further performed an enhanced sampling simulation to analyze the ring formation process at $\sim$ 10 GPa and 1400 K.
 The simulation box has one “C4N2-2” molecule and 32 water molecules, as shown in Fig.
@@
 Geochemical evidence of biotic activity suggests that the emergence of life can be traced back to as early as 3.8 billion years ago mojzsis1996evidence .
-The modern style of plate tectonics may have started about 3.2 billion years ago brenner2020paleomagnetic , but some form of plate tectonics may have already been active around 4.0 billion years ago.
-and these tectonic activities might carry water, carbon, and other volatile substances down into the mantle korenaga2013initiation ; HOLDER2024 ; blichert2008hafnium ; hopkins2008low ; huang2022barium .
-Because of reducing environments in early Earth’s mantle aulbach2016evidence , oxidized carbon and nitrogen can be reduced to CO and NH3/NH+${}_{4}^{+}$, respectively holm2009reduction (see Figure S16), and they may further synthesize to become key building blocks of life, which may reach the surface of early Earth.
+The modern style of plate tectonics may have started about 3.2 billion years ago brenner2020paleomagnetic , but some form of plate tectonics may have already been active around 4.0 billion years ago, and these tectonic activities might carry water, carbon, and other volatile substances down into the mantle korenaga2013initiation ; HOLDER2024 ; blichert2008hafnium ; hopkins2008low ; huang2022barium .
+Because of reducing environments in early Earth’s mantle aulbach2016evidence , oxidized carbon and nitrogen can be reduced to CO and NH3/NH+${}_{4}^{+}$, respectively (see Figure S16)holm2009reduction , and they may further synthesize to become key building blocks of life, which may reach the surface of early Earth.
 Our findings also have important implications for organic molecules in interstellar space.
@@
 It is unclear how those complex organic species are generated.
-Our current study suggests that the extreme pressure and temperature inside giant planets and their moons neri2020carbonaceous ; nettelmann2016uranus , like Jupiter and Saturn, may help to synthesize those complex organic molecules postberg2018macromolecular ; waite2009liquid ; tosi2024salts .
+Our study suggests that the extreme P and T inside giant planets and their moons neri2020carbonaceous ; nettelmann2016uranus , like Jupiter and Saturn, may help to synthesize those complex organic molecules postberg2018macromolecular .
+Earth’s or other giant planets’ interiors also provide a protective environment that filters out harmful radiation and buffers against drastic physico-chemical variations.
 III Conclusions
@@
 Three corresponding P-T systems were considered, i.e., 10 GPa - 1000 K, 10 GPa - 1400 K, and 13 GPa - 1400 K.
-The simulation results have revealed the potential formation of substantial quantities of organic matter without any catalysts, with the composition and abundance of CN-containing species being profoundly influenced by the specific P-T conditions.
+The simulation results revealed the potential formation of substantial quantities of organic matter without any catalysts, with the composition and abundance of CN-containing species being profoundly influenced by the specific P-T conditions.
 Notably, formamide, recognized as a crucial precursor in the production of biomolecules, is observed to readily form under extreme P-T conditions, and its formation pathways are compared across different P-T conditions.
-Furthermore, free energy calculations have confirmed that the 10 GPa - 1400 K condition is particularly favorable for the generation and preservation of CN-containing compounds.
+Furthermore, free energy calculations confirmed that the 10 GPa - 1400 K condition is particularly favorable for the generation and preservation of CN-containing compounds.
 Using the molecules formed, we have further explored the formation process of large biomolecules that are directly related to the building blocks of life.
 Remarkably, glycine, ribose, urea, and uracil-like molecules are successfully synthesized under extreme conditions in our AIMD simulations, thereby providing initial insights into the formation of organic compounds critical to the emergence of life.
-More importantly, different from that the six-membered ring of ribose is the majority in solutions at ambient conditions, it is found that formation of five-membered ring form of ribose is more favored at extreme P-T conditions, which could provide the necessary components for the generation of RNA molecules.
-Moreover, these compounds have the potential to be transported to Earth’s surface through geological activities, such as mantle convection, consequently contributing to the origins of life.
+Particularly, different from that the six-membered ring form of ribose is the majority in solutions at ambient conditions, it was found that the formation of the five-membered ring form of ribose is more favored at extreme P-T conditions, which provides the necessary component for the generation of RNA molecules.
+Moreover, these compounds have the potential to be transported to Earth’s surface through geological activities, such as mantle convection, consequently contributing to the origin of life.
 The complex organic species found in interstellar space may be also related to the chemical processes inside giant planets and their moons at extreme P-T conditions.
-Our comprehensive studies present a novel perspective extending the cradle of life planetary interiors.
+Our comprehensive studies present a novel perspective extending the cradle of life in planetary interiors.
 The findings underscore the profound impact of extreme P-T conditions on the formation of organic matter and provide valuable insights into the mechanisms underlying the emergence of life within planetary bodies.
@@
 The simulation boxes have periodic boundary conditions.
-We first performed NPT simulations to achieve the desired pressures, configuring the plane-wave energy cutoff to 85 Ry and the energy cutoff for stress confinement potential to 70 Ry.
-We applied the simulation box sizes (Table S1) obtained from NPT simulations in the following NVT simulations for production runs, where the plane-wave energy cutoff was 65 Ry.
-During NVT simulations, we selected random time points to carry out independent simulations with energy cutoff of 85 Ry to monitor pressure (Figure S2).
+We first performed NPT simulations to achieve the desired pressures, configuring the plane-wave kinetic energy cutoff to 85 Ry and the cutoff for stress confinement potential to 70 Ry focher1994structural .
+We applied the simulation box sizes (Table S1) obtained from NPT simulations in the following NVT simulations for production runs, where the plane-wave kinetic energy cutoff was 65 Ry.
+During NVT simulations, we selected random time points to carry out independent simulations with the cutoff of 85 Ry to monitor pressure (Figure S2).
 Our total AIMD simulation time exceeds 2.5 ns.
@@
 To compute the free-energy landscape, we employed the adaptive biasing force method (ABF) darve2008adaptive as implemented in the software package SSAGES sidky2018ssages coupled with Qbox code.
-Unlike other advanced sampling methods that introduce biases into the energy, this method aims to flatten the generalized force.
+Unlike other advanced sampling methods that introduce bias potentials, this method aims to flatten the generalized force.
 The running average of the force in the $k$th bin along the reaction coordinate $\xi$ is calculated as:
@@
 $F_{i}(t_{i}^{k})$ $=\frac{d}{dt}\left(M_{\xi}\frac{d\xi}{dt}\right)\bigg{|}_{t_{i}^{k}}$ (2)
-where $\nu\left({{N}_{step}},k\right)$ is the number of samples collected in the $k$th bin after $N_{step}$ steps, ${{F}_{i}}\left(t_{i}^{k}\right)$ is the $i$th force sample at the time $t_{i}^{k}$, M is the mass matrix.
+where $\nu\left({{N}_{step}},k\right)$ is the number of samples collected in the $k$th bin after $N_{step}$ steps, ${{F}_{i}}\left(t_{i}^{k}\right)$ is the $i$th force sample at the time $t_{i}^{k}$, and M is the mass matrix.
 The bias force -$\text{-}\left(\nabla\xi\right){{F}_{\xi}}\left({{N}_{step,}}k\right)$ is added to the calculation until the free energy landscape is flat.
@@
 […]
-Supporting information
-Details of the simulations, the percentage of C-N containing species over time, the hydrocarbon products, the free energy calculations, other strategies for the formation of ribose and uracil, and their ring structures, compositions of fluids at different oxygen fugacity.
 Acknowledgements

### v4

@@
-Present address: ]Lehrstuhl für Theoretische Chemie, Ruhr-Universität Bochum, 44780 Bochum, Germany
+Synthesis and stability of biomolecules in C-H-O-N fluids under Earth’s upper mantle conditions
 Abstract
@@
 10 - 13 GPa, 1000 - 1400 K).
-Contrary to the previous assumptions that larger organic molecules might readily disintegrate in aqueous solutions at extreme P-T conditions, we found that many organic compounds formed without any catalysts and persisted in C-H-O-N fluids under these extreme conditions, including glycine, ribose, urea, and uracil-like molecules.
+Contrary to the previous assumptions that large organic molecules might readily disintegrate in aqueous solutions at extreme P-T conditions, we found that many organic compounds formed without any catalysts and persisted in C-H-O-N fluids under these extreme conditions, including glycine, ribose, urea, and uracil-like molecules.
 Particularly, our free energy calculations showed that the C-N bond is thermodynamically stable at 10 GPa and 1400 K.
-Moreover, while the pyranose (six-membered-ring) form of ribose is more stable than the furanose (five-membered-ring) form at ambient conditions, we observed the predominant formation of the five-membered-ring form of ribose at extreme conditions, which is consistent with the exclusive incorporation of $\beta$-D-ribofuranose in RNA.
-We have uncovered a previously unexplored pathway through which the building blocks of biomolecules may have originated in early Earth and other planets.
-Our findings contribute to an evolving understanding of the fundamental conditions necessary for life to arise.
+Moreover, while the pyranose (six-membered-ring) form of ribose is more stable than the furanose (five-membered-ring) form at ambient conditions, we found that the formation of the five-membered-ring form of ribose is thermodynamically more favored at extreme conditions, which is consistent with the exclusive incorporation of $\beta$-D-ribofuranose in RNA.
+We have uncovered a previously unexplored pathway through which the crucial biomolecules could be abiotically synthesized from geofluids in the deep interior of Earth and other planets and these formed biomolecules could potentially contribute to the early stage of the emergency of life.
 I Introduction
@@
 Those deep sea vents were considered to have suitable conditions for the origin of life martin2008hydrothermal , but previous studies also suggested that the high temperatures in the vents could readily degrade crucial biomolecules in aqueous solutions miller1988submarine .
-Shock recovery experiments successfully synthesized various biomolecules, where the pressure can reach tens of gigapascals and the temperature can rise to thousands of kelvins before releasing back to ambient conditions Sekine2024 ; however, due to the dynamic nature of shock waves, samples do not reach thermodynamic equilibrium states, so the shock experiments are usually conducted to mimic the conditions of meteorite impacts on Earth furukawa2009biomolecule .
+Shock recovery experiments successfully synthesized various biomolecules, where the pressure can reach tens of gigapascals and the temperature can rise to thousands of kelvins before releasing back to ambient conditions Sekine2024 .
+However, due to the dynamic nature of shock waves, samples do not reach thermodynamic equilibrium states, so the shock experiments are usually conducted to mimic the conditions of meteorite impacts on Earth furukawa2009biomolecule .
 It was commonly assumed that large biomolecules could not persist at both high hydrostatic pressures and high temperatures as found in Earth’s upper mantle, where pressures can reach up to $\sim$13 GPa, accompanied by high temperatures of $\sim$1700 K.
 In recent years, Otake et al.
-studied the stability of amino acids and their oligomerization at 1.0–5.5 GPa and 453–673 K, and found that high-pressure conditions inhibited the decomposition of amino acids and small amino acids could be oligomerized up to pentamers; however, their experiments were conducted at the water-poor condition otake2011stability .
+studied the stability of amino acids and their oligomerization at 1.0–5.5 GPa and 453–673 K, and found that high-pressure conditions inhibited the decomposition of amino acids and small amino acids could be oligomerized up to pentamers.
+However, their experiments were conducted at the water-poor condition otake2011stability .
 In aqueous solutions, there seems a strong thermodynamic drive towards the formation of amino acids dimers above 1.5 GPa and temperature above 423 K robinson2021quantifying , in agreement with some experimental observations pedreira2019spontaneous .
@@
 These findings have challenged the previous assumptions and shed light on the potential existence of larger organic molecules in Earth’s or other planets’ subsurface or even deeper.
-It is very challenging to study the origin of life in laboratory settings due to the extreme environments and subsequent analyses.
+It is very challenging to study the building blocks of biomolecules for life in laboratory settings due to the extreme environments and subsequent analyses.
 Atomistic simulations based on first principles have been applied to study prebiotic chemistry, offering significant molecular-level insights perez2020prebiotic .
 Saitta et al.
-applied ab initio molecule dynamics (AIMD) and metadynamics simulations to study the Miller-Urey experiment, and found unexpected reaction intermediates including formic acid and formamide saitta2014miller .
+applied ab initio molecular dynamics (AIMD) and metadynamics simulations to study the Miller-Urey experiment, and found unexpected reaction intermediates including formic acid and formamide saitta2014miller .
 Goldman et al.
@@
 Despite the numerous previous studies, the exploration of prebiotic chemistry at both high hydrostatic pressures and high temperatures largely remains an uncharted territory.
-Here, we applied extensive AIMD simulations to study the C-H-O-N fluids at 10–13 GPa and 1000–1400 K, the P-T conditions as found in the transition zone of Earth’s upper mantle.
+Here, we applied extensive AIMD simulations to study the reactions of inorganic C-H-O-N geofluids at 10–13 GPa and 1000–1400 K, the P-T conditions as found in the transition zone of Earth’s upper mantle.
 We discovered that a large range of organic molecules directly formed without any catalysts at extreme P-T conditions, including formamide, glycine, ribose, urea, and uracil-like molecules.
-Our free energy calculation suggested that the C-N bond is thermodynamically stable at 10 GPa and 1400 K.
-Our study provides valuable insights into a new potential avenue for the formation of building blocks of life and potentially for the origin of life within Earth’s and other planets’ interior, expanding our understanding of the conditions necessary for life to emerge.
+The formation pathways are also detailed discussed.
+Our free energy calculations using AIMD simulations suggest that the C-N bond is thermodynamically stable at 10 GPa and 1400 K.
+Our study suggests a new potential avenue for the abiotic formation of biomolecules in the deep interior of Earth and other planets, further expanding the applicable places of “primordial soup” theory.
 II Results and Discussion
-The simulation box initially contained 15 molecules of each of the following species: H2O, H2, CO, and NH3 (Figure S1).
+The simulation box initially contained 15 molecules of each of the following species: H2O, H2, CO, and NH3 (Fig.
+S1).
 Those molecules are typical volatile compounds found in C-H-O-N fluids in reducing environments of Earth’s deep interior zhang2009model ; manning2013chemistry , and are also abundant in interstellar space guelin2022organic .
@@
 1 shows the chemical structures of the representative species.
-Species with more than five carbon atoms were observed; however they are excluded at this stage due to their high instability and extremely short lifetime, typically lasting for only a few femtoseconds.
-The percents of these CN-containing species as a function of time at three different P-T conditions are shown in Figure S3, which suggest that the concentrations of these species gradually stabilize over time.
+Species with more than five carbon atoms were observed, however they are excluded at this stage due to their high instability and extremely short lifetime, typically lasting for only a few femtoseconds.
+The percents of these CN-containing species as a function of time at three different P-T conditions are shown in Fig.
+S3, which suggest that the concentrations of these species gradually stabilize over time.
 At 10 GPa and 1000 K, most of the organic products belong to three main categories, as shown in Fig.
@@
 Fig.
-3 (a) shows the different formation pathways at three P-T conditions, and detailed formation processes are displayed in Figure S4.
+3 (a) shows the different formation pathways at three P-T conditions, and detailed formation processes are displayed in Fig.
+S4.
 At 10 GPa and 1000 K, the formation reaction begins with the dissociation of water molecules, generating OH- and H+ ions in the aqueous solution.
@@
 For the generation of biomolecules in living matter, the formation of C-N bonds is the key step.
-We further applied the advanced sampling method (see details in the Methods section) to calculate the free energy profile of the C-N bond formation at three P-T conditions in Fig.
+We further applied the advanced sampling method (see details in the Computational Methodology section) to calculate the free energy profile of the C-N bond formation at three P-T conditions in Fig.
 3(b).
@@
 The free energy profiles suggest that the energy barriers of the C-N bond dissociation are 17.57 kcal/mol and 16.40 kcal/mol, at 10 GPa, 1000 K and 13 GPa, 1400 K, respectively.
-As a comparison, at 10 GPa and 1400 K, the free energy increases upon dissociation of the C-N bond (1.31 Å), and the energy barrier is 59.04 kcal/mol, indicating that the C-N bond is thermodynamically and kinetically stable at this P-T condition.
+As a comparison, at 10 GPa and 1400 K, the free energy increases upon dissociation of the C-N bond (1.31 Å), and the energy barrier is 49.04 kcal/mol, indicating that the C-N bond is thermodynamically and kinetically stable at this P-T condition.
 The free energy result agrees with the data in Fig.
@@
 4 (a) illustrates the process of forming $\alpha$-hydroxyglycine from formamide.
-A proton attacked the formamide molecule, leading to the creation of a protonated formamide molecule, also known as an aminohydroxymethyl radical.
-Then, a CO2 molecule reacts with the protonated formamide, resulting in the “-N-C-C-” backbone through forming a C-C bond.
+A proton attacks the formamide molecule, leading to the creation of a protonated formamide molecule.
+Then, the protonated formamide reacts with a large CO2-containing molecule, resulting in the “-N-C-C-” backbone through forming a C-C bond.
 Finally, another proton is captured by a “-COO” group, ultimately yielding the $\alpha$-hydroxyglycine molecule.
@@
 Considering that the CN-backbone of the “C2N” species at 13 GPa - 1400 K is “-C-N-C-”, different from the CN-backbone in the glycine molecule, we did not consider this P-T condition.
-We can see that the free energy minima is at 1.07 Å for both 10 GPa, 1000 K and 10 GPa, 1400 K, and the energy barriers to break the C-H bond are 21.91 kcal/mol and 22.04 kcal/mol, respectively, which are very similar.
+We can see that the free energy minima is at 1.07 Å for both 10 GPa, 1000 K and 10 GPa, 1400 K, and the energy barriers to break the C-H bond are 21.91 kcal/mol and 22.14 kcal/mol, respectively, which are very similar.
 After the C-H bond dissociation, the free energy increases more at 10 GPa and 1400 K (6.76 kcal/mol) than at 10 GPa and 1000 K (2.63 kcal/mol), indicating that the C-H bond is more favored at the former P-T condition.
@@
 Initially, two $\alpha$-hydroxyglycine molecules combine to form one molecule, but later dissociate to dehydroglycine and glyoxylic acid after $\sim$28 ps.
-As the simulation goes on, the C=N double bond in the dehydroglycine molecule receives two protons and finally dehydroglycine becomes glycine at $\sim$ 38 ps.
+As the simulation goes on, the dehydroglycine will be reduced by H2 and finally it becomes glycine at $\sim$ 38 ps.
 Both proteins and DNA are widely recognized as the important biomolecules for building blocks of life.
@@
 The formose reaction is a well-known prebiotic pathway for synthesizing ribose from formaldehyde (CH2O) breslow1959mechanism .
-Our simulations have generated a few relevant hydrocarbon products (shown in Figures S5 and S6), including formaldehyde molecule (number 1 in Figure S5), which was later chosen as the primary reactant.
-Additionally, to facilitate the synthesis process, the reactants also included a “C2” molecule, (Z)-ethene-1,2-diol (C2H4O2, number 7 in Figure S5).
-Initially, we constructed a simulation box consisting of 10 (Z)-ethene-1,2-diol molecules and 5 formaldehyde molecules (Figure S7).
+Our simulations have generated a few relevant hydrocarbon products (shown in Figs.
+S5 and S6), including formaldehyde molecule (number 1 in Fig.
+S5), which was later chosen as the primary reactant.
+Additionally, to facilitate the synthesis process, the reactants also included a “C2” molecule, (Z)-ethene-1,2-diol (C2H4O2, number 7 in Fig.
+S5).
+Initially, we constructed a simulation box consisting of 10 (Z)-ethene-1,2-diol molecules and 5 formaldehyde molecules (Fig.
+S7).
 After $\sim$76 ps, a “C5” molecule was formed at $\sim$10 GPa and 1400 K.
@@
 By measuring the distances of C1-C1 and C2-C2 , we confirmed that they are consistent to the distances observed in the free energy landscape.
-We also randomly selected another initial configuration to confirm that the ribose molecule is thermodynamically stable, as shown in Figure S8.
+Furthermore, another initial configuration for the free energy calculation is also presented in Fig.
+S8, the obtained free energy landscape still shows that the ribose molecule is thermodynamically stable.
 Previous studies showed that the cyclic forms of ribose are abundant in aqueous solutions at room temperature drew199813c .
@@
 6 (a).
-Previous experimental studies on the origin of life have suggested that formamide is a crucial molecule for the prebiotic formation of nucleobases, in the presence of mineral or metal oxide catalysts saladino2012formamide , heatingenchev2021chemical , and high-energy free radicals ferus2015high ; ferus2014high , so here we also first chose formamide as a reactant.
-However, after about 144 ps, we did not observe the formation of the “C4N2” molecule at 10 GPa and 1400 K, suggesting that it may be difficult for small molecules to directly synthesize uracil in the AIMD simulations (Figure S9).
+Previous experimental studies on the origin of life suggested that formamide is a crucial molecule for the prebiotic formation of nucleobases, in the presence of mineral or metal oxide catalysts saladino2012formamide , heatingenchev2021chemical , and high-energy free radicals ferus2015high ; ferus2014high , so here we also first chose formamide as a reactant.
+However, after about 144 ps, we did not observe the formation of the “C4N2” molecule at 10 GPa and 1400 K, suggesting that it may be difficult for small molecules to directly synthesize uracil in the AIMD simulations (Fig.
+S9).
 Later, we chose the chemical reactants that have closer structural similarities to the “C4N2” molecule.
 The reactant molecules, including “C4N”, “C3N2”, “C2N2”, and “C3”, were generated from our first-stage AIMD simulations.
-A recent study suggested that uracil could be synthesized from the reactions between urea and “C3” molecules choe2022mechanism , so we specifically selected the C3H4O4 molecule (number 10 in Figure S5) and the urea molecule (CH4N2O, number 11 in Fig.
+A recent study suggested that uracil could be synthesized from the reactions between urea and “C3” molecules choe2022mechanism , so we specifically selected the C3H4O4 molecule (number 10 in Fig.
+S5) and the urea molecule (CH4N2O, number 11 in Fig.
 1) as reactants.
-The simulation results are presented in Figure S10.
+The simulation results are presented in Fig.
+S10.
 However, regardless of the inclusion of water molecules or the application of different pressure conditions, we did not observe any “C4N2” molecules with the CN-backbone of the open-chain form of uracil.
-Then we tried to find another pathways and conducted six more AIMD simulations to investigate the possibility of forming uracil through reactions involving above-mentioned molecules at various pressures and 1400 K (see Figures S11, S12 and S14).
+Then we tried to find another pathways and conducted six more AIMD simulations to investigate the possibility of forming uracil through reactions involving above-mentioned molecules at various pressures and 1400 K (see Figs.
+S11, S12 and S14).
 […]
@@
 Later we chose allophanic acid (C2H4N2O3, number 13 in Fig.
-1 ) and (Z)-ethene-1,2-diol (C2H4O2, number 7 in Figure S5) as reactants, and built a new simulation box containing 5 C2H4N2O3, 5 C2H4O2, and 10 H2, as shown in Fig.6 (c).
-At $\sim$ 10 GPa and 1400 K, a “C5N2” molecule, 3-(3-carboxyureido)-2,3-dihydroxypropanoic acid, was formed at around 28.8 ps, This molecule has the same CN-backbone as the open-chain form of thymine (T).
+1 ) and (Z)-ethene-1,2-diol (C2H4O2, number 7 in Fig.
+S5) as reactants, and built a new simulation box containing 5 C2H4N2O3, 5 C2H4O2, and 10 H2, as shown in Fig.6 (c).
+At $\sim$ 10 GPa and 1400 K, a “C5N2” molecule, 3-(3-carboxyureido)-2,3-dihydroxypropanoic acid, was formed at around 28.8 ps.
+This molecule has the same CN-backbone as the open-chain form of thymine (T).
 As the simulation continued, a “C4N2” molecule, 2,3-dihydroxy-3-ureidopropanoic acid, labeled as “C4N2-2”, was formed, which has the same CN-backbone as the open-chain form of uracil.
-The formation process of these two kinds of molecules are shown in Figure S15.
 We further performed an enhanced sampling simulation to analyze the ring formation process at $\sim$ 10 GPa and 1400 K.
@@
 6 (c) shows that the cyclic form is more stable than the open-chain form, indicating that the uracil-like molecule is possible to form in geological C-H-O-N fluids at extreme P-T conditions.
+The formation process and reaction mechanism of the uracil-like molecule are shown in Fig.
+S15.
 The uracil-like molecule can then serve as a precursor for the synthesis of other three types of nucleobases, providing the necessary components for forming RNA molecules.
 Geochemical evidence of biotic activity suggests that the emergence of life can be traced back to as early as 3.8 billion years ago mojzsis1996evidence .
-The modern style of plate tectonics may have started about 3.2 billion years ago brenner2020paleomagnetic , but some form of plate tectonics may have already been active around 4.0 billion years ago, and these tectonic activities might carry water, carbon, and other volatile substances down into the mantle korenaga2013initiation ; HOLDER2024 ; blichert2008hafnium ; hopkins2008low ; huang2022barium .
-Because of reducing environments in early Earth’s mantle aulbach2016evidence , oxidized carbon and nitrogen can be reduced to CO and NH3/NH+${}_{4}^{+}$, respectively (see Figure S16)holm2009reduction , and they may further synthesize to become key building blocks of life, which may reach the surface of early Earth.
-Our findings also have important implications for organic molecules in interstellar space.
+The modern style of plate tectonics may have started about 3.2 billion years ago brenner2020paleomagnetic , but some forms of plate tectonics may have already been active around 4.0 billion years ago, and these tectonic activities might carry water, carbon, and other volatile substances down into the mantle korenaga2013initiation ; HOLDER2024 ; blichert2008hafnium ; hopkins2008low ; huang2022barium .
+Because of reducing environments in early Earth’s mantle aulbach2016evidence , oxidized carbon and nitrogen can be reduced to CO and NH3/NH+${}_{4}^{+}$, respectively (see Fig.
+S16)holm2009reduction , and they may further synthesize to become key building blocks of life, which may reach the surface of early Earth.
+Our findings also have important implications for the synthesis of organic molecules in interstellar space.
 The reactant molecules H2O, H2, CO and NH3 largely exist in interstellar clouds and circumstellar shells, together with more than 200 complex organic species, such as aldehydes, alcohols, acids, amines and carboxamides guelin2022organic .
-It is unclear how those complex organic species are generated.
-Our study suggests that the extreme P and T inside giant planets and their moons neri2020carbonaceous ; nettelmann2016uranus , like Jupiter and Saturn, may help to synthesize those complex organic molecules postberg2018macromolecular .
+The formation mechanism of those complex organic species may be intricate.
+Our study suggests that the extreme P and T inside giant planets and their moons neri2020carbonaceous ; nettelmann2016uranus , like Jupiter and Saturn, could be considered as one important condition to help to synthesize some of those complex organic molecules postberg2018macromolecular .
 Earth’s or other giant planets’ interiors also provide a protective environment that filters out harmful radiation and buffers against drastic physico-chemical variations.
+Although our simulations were conducted under extreme P-T conditions, our study may also have implications for experiments at lower pressures and temperatures.
+We found that high pressures can facilitate the synthesis and stabilization of large biomolecules, aligning with some experimental studies conducted at lower pressures otake2011stability ; ohara2007pressure ; furukawa2012abiotic ; pedreira2019spontaneous .
+It’s important to note that some experiments were performed under water-poor conditions otake2011stability ; ohara2007pressure ; furukawa2012abiotic , whereas we discovered that supercritical water plays a crucial role in the formation of biomolecules.
 III Conclusions
@@
 Furthermore, free energy calculations confirmed that the 10 GPa - 1400 K condition is particularly favorable for the generation and preservation of CN-containing compounds.
-Using the molecules formed, we have further explored the formation process of large biomolecules that are directly related to the building blocks of life.
+Using the molecules formed, we have further explored the formation pathways of large biomolecules that are directly related to the building blocks of life.
 Remarkably, glycine, ribose, urea, and uracil-like molecules are successfully synthesized under extreme conditions in our AIMD simulations, thereby providing initial insights into the formation of organic compounds critical to the emergence of life.
 Particularly, different from that the six-membered ring form of ribose is the majority in solutions at ambient conditions, it was found that the formation of the five-membered ring form of ribose is more favored at extreme P-T conditions, which provides the necessary component for the generation of RNA molecules.
-Moreover, these compounds have the potential to be transported to Earth’s surface through geological activities, such as mantle convection, consequently contributing to the origin of life.
+Moreover, these compounds also have the potential to be transported to Earth’s surface through geological activities, such as mantle convection, consequently contributing to the origin of life.
 The complex organic species found in interstellar space may be also related to the chemical processes inside giant planets and their moons at extreme P-T conditions.
-Our comprehensive studies present a novel perspective extending the cradle of life in planetary interiors.
-The findings underscore the profound impact of extreme P-T conditions on the formation of organic matter and provide valuable insights into the mechanisms underlying the emergence of life within planetary bodies.
-Methods
+Our comprehensive studies underscore the profound impact of extreme P-T conditions on the abiotic synthesis of organic matter and provide valuable insights into the mechanisms underlying the formation of crucial biomolecules for building blocks of life, which presents a novel perspective extending the cradle of life in planetary interiors.
+Computational Methodology
 III.1 Ab initio molecular dynamics simulations (AIMD)
@@
 We applied the simulation box sizes (Table S1) obtained from NPT simulations in the following NVT simulations for production runs, where the plane-wave kinetic energy cutoff was 65 Ry.
-During NVT simulations, we selected random time points to carry out independent simulations with the cutoff of 85 Ry to monitor pressure (Figure S2).
+During NVT simulations, we selected random time points to carry out independent simulations with the cutoff of 85 Ry to monitor pressure (Fig.
+S2).
+For each P-T point, the NVT simulation time exceeds 400 ps (Fig.
+S3).
 Our total AIMD simulation time exceeds 2.5 ns.
+In our AIMD simulations, we calculated the forces on the nuclei using quantum mechanics, while treating nuclei including protons as classical particles.
+If we consider quantum nuclear effects, we anticipate an increase in the proton transfer rate, resulting in more active reactions involving proton transfer.
+However, at high temperatures, the significance of quantum effects may diminish compared to ambient conditions ceriotti2016nuclear .
 Free-energy calculations by coupling AIMD simulations with enhanced sampling
@@
 […]
+In our simulations, the calculated free energies are converged within 0.2 kcal/mol in the last 10 ps. The test of convergence of free energy calculations is shown in Fig. S17.
+Supporting information
+Details of the simulations, Mole percents of C-N containing species over time, hydrocarbon products, free energy calculations, possible synthesis of ribose and uracil and their ring structures, compositions of fluids as a function of oxygen fugacity, test of convergence of free energy calculations.
 Acknowledgements
-This work was supported by the Croucher Foundation through the Croucher Innovation Award, Hong Kong Research Grants Council (Projects GRF-16301723, GRF-16306621, and C6021-19EF), National Natural Science Foundation of China through the Excellent Young Scientists Fund (22022310), and the Hetao Shenzhen/Hong Kong Innovation and Technology Cooperation (HZQB-KCZYB-2020083).
+This work was supported by the Croucher Foundation through the Croucher Innovation Award, Hong Kong Research Grants Council (Projects GRF-16301723, GRF-16306621, and C6021-19EF), National Natural Science Foundation of China through the Excellent Young Scientists Fund (22022310).
 R.T.